### v2

# DCoM: Active Learning for All Learners

Abstract
Deep Active Learning (AL) techniques can be effective in reducing annotation costs for training deep models. However, their effectiveness in low- and high-budget scenarios seems to require different strategies, and achieving optimal results across varying budget scenarios remains a challenge. In this study, we introduce Dynamic Coverage & Margin mix (DCoM), a novel active learning approach designed to bridge this gap. Unlike existing strategies, DCoM dynamically adjusts its strategy, considering the competence of the current model. Through theoretical analysis and empirical evaluations on diverse datasets, including challenging computer vision tasks, we demonstrate DCoM’s ability to overcome the cold start problem and consistently improve results across different budgetary constraints. Thus DCoM achieves state-of-the-art performance in both low- and high-budget regimes.
1 Introduction
Deep Learning (DL) algorithms require a lot of data to achieve optimal results. In certain scenarios, there is an abundance of unlabeled data but limited capacity for labeling it. In fields such as medical imaging, an invaluable resource — doctors themselves — serves as a costly oracle. Active learning algorithms aim to address this challenge by reducing the burden of labeling and making it more effective. Unlike traditional supervised learning frameworks, Active Learning (AL) can affect the construction of the labeled set, possibly by leveraging knowledge about the current learner. Accordingly, the initial goal of AL is to select $q$ examples to be annotated, where $q$ represents the number of examples that can be sent to the oracle. AL has already demonstrated tangible contributions across various domains such as computer vision tasks [40], NLP [31, 20] and medical imaging [10]. These examples highlight the importance of advancing AL to achieve even greater impact.
The choice of an AL strategy depends on both the learner’s inductive biases and the nature of the problem. Recent research suggests that the optimal active learning strategy varies with the budget size, where budget refers to the size of the training set. When the budget is large, methods based on uncertainty and diversity sampling are most effective. Conversely, when the budget is small, methods centered on typicality and diversity are more appropriate. However, there is no single active learning strategy that is suitable for all budget regimes. Our study aims to address this challenge by dynamic selection of the best examples based on the current state of the learner and the current budget.
As briefly reviewed below, recent research often categorizes active sampling methods according to the budget regime. In contrast, our study takes a different approach by recognizing the need to adapt sample selection methods based on the learner’s competence. We aim to develop a single algorithm that dynamically adjusts its sampling strategy in response to the learner’s evolving abilities. This transition from budget to learner’s competence represents a significant advancement in active learning research, as it allows us to move beyond rigid budget allocations and towards a more adaptive approach that optimizes performance based on the learner’s abilities.
In this work, we introduce a novel method called Dynamic Coverage & Margin mix (DCoM), which dynamically adjusts its selection strategy to deliver optimal results. We begin by examining a simplified theoretical framework (Section 2), where we integrate a typicality test with an uncertainty test based on the budget and its coverage. Inspired by the insights gained from this analysis, we introduce DCoM (Section 3), which combines these characteristics. DCoM aims to offer a flexible solution for any budget by selecting the most appropriate examples that are most suitable to the current budget. We validate DCoM through an extensive empirical study using various vision datasets (Section 4). Our findings indicate that DCoM effectively achieves superior performance across all budget ranges.
Relation to prior art. Over the past years, active learning has been an active area of research [34, 29, 7, 32]. The realm of Active Learning (AL) is segmented into three problem settings: membership query synthesis, stream-based selective sampling, and pool-based active learning [29]. The majority of current AL research adheres to the pool-based AL setting, where samples are drawn from a vast pool of unlabeled data and annotations are solicited from an oracle. In this setting, AL algorithms operate within a given budget ($\mathit{b}$), selecting a subset of unlabeled examples to send to an oracle for labeling. This selection process aims to identify examples that will maximize the model’s performance. This process may be repeated iteratively, gradually increasing budget $b$.
In discussions of pool-based AL, two key characteristics that play a crucial role during the active sampling process are uncertainty and diversity. Some AL algorithms focus solely on uncertainty sampling [23, 30, 4, 43, 18, 12, 27, 37, 11], while others prioritize diversity sampling [28, 19, 39, 15]. Other algorithms combine both uncertainty and diversity in their approaches [38, 2, 36, 21, 5, 1].
Over time, the cold start problem emerged, highlighting the challenges faced by AL algorithms when operating with a small budget. It seems that uncertainty and diversity do not capture this difficulty. Chen et al. [2], for example, describe the cold start problem as a combination of two issues: unbalanced sampling across different classes (which relates to diversity) and unbalanced sampling within each class, where uncertain examples are favored over certain ones at the outset.
In low-budget scenarios, typical examples often offer the best outcomes for the model [2, 39, 15, 14]. This implies that in such contexts, AL algorithms should aim to maximize both certainty and diversity. Conversely, in high-budget scenarios, the emphasis shifts towards maximizing uncertainty while maintaining diversity. [14] proposes a method to select the most suitable approach based on the current state. This method involves comparing multiple active learning strategies during runtime, rather than recommending a single universal approach.
Summary of contribution. (i) Expand a theoretical framework to analyze Active Learning (AL) strategies in embedding spaces. (ii) Introduce DCoM, a multi-budget strategy that significantly outperforms previous methods in low and medium budget scenarios and matches their performance in high budget scenarios. (iii) Initiates a transition from a budget-based approach to the learner’s competence, introducing coverage as a measure designed to predict this competence.
2 Theoretical framework
The framework of active learning adopted here can be formalized as follows: Let ${\mathbb{X}}$ denote the input domain, ${\mathbb{Y}}$ the target domain, and $\mathcal{L}:{\mathbb{X}}\to{\mathbb{Y}}$ the target learner. Before seeing any labels, $\mathcal{L}$ may be a random feasible hypothesis, or it may be initialized by either transfer learning or self training.
Active learning may involve a single step, or an iterative process with repeated active learning steps. In each step of active a learning, the learner actively seeks labels by choosing a set of unlabeled points as a query set, to be labeled by an oracle/teacher. Subsequently, this set of labeled points is added to the learner’s supervised training set, and the learner is retrained or fine-tuned.
As discussed in the introduction, it has been shown that different active learning strategies are suitable for different budgets $b$. Henceforth, $b$ refers to the number of labeled examples known to the learner at the beginning of an active learning step. With a low budget $b$, strategies that do not rely on the outcome of the learner are most suitable. When $b$ is high, it is beneficial to take into consideration the confidence of the learner when choosing an effective query set. What makes a budget high or low was left vague in previous analysis, as it clearly depends on the specific application and dataset.
To achieve an effective active learning protocol suitable for all learners, irrespective of budget, we aim to devise a universal objective function, whose minimization is used to select the query set. Next, we discuss its envisioned form for selecting the query set at the beginning of each AL step.
We begin our discussion by proposing to replace the notion of budget with the notion of competence: a network is competent if its generalization score is relatively high, and vice versa. This proposal is more formally justified in Section 2.3. We then rephrase the intuition stated above as follows: When the learner’s competence is low, the objective function should prioritize typicality and diversity of selected queries (see discussion in the introduction) irrespective of the learner’s predictions. When the learner’s competence is high, its uncertainty in prediction should be given high priority.
More formally, let ${\mathcal{O}(x)}$ denote the desired objective function for query selection, where $x\in{\mathbb{X}}$ is unlabeled. Let ${{\mathcal{O}_{low}}(x)}$ and ${{\mathcal{O}_{high}}(x)}$ denote the objective functions suitable for a learner with low competence and high competence respectively. Let ${S_{\mathcal{L}}}$ denote a score, which captures the competence of learner $\mathcal{L}$. The proposed objective function can now be written as follows $\forall x\in{\mathbb{X}}$:
${\mathcal{O}}(x)=(1-{S_{\mathcal{L}}})\cdot{{{\mathcal{O}_{low}}}(x)}+{S_{\mathcal{L}}}\cdot{{{\mathcal{O}_{high}}}(x)}$ (1)
In Section 2.2 we discuss the design of ${\mathcal{O}_{low}}(x)$. In order to tackle the challenge of active learning when the competence of $\mathcal{L}$ is low, we utilize the point coverage framework introduced in [39]. This approach relies on self-supervised data representation, which is blind to the learner’s performance. We refine their analysis and obtain an improved generalization bound for the 1-Nearest Neighbor (1-NN) classification model that depends on the local geometry of the data. Beforehand, in Section 2.1 we discuss the necessary preliminaries required to introduce the coverage approach. In Section 2.3 we discuss our choice of ${S_{\mathcal{L}}}$, and connect it to the notion of budget discussed in previous work.
We select ${\mathcal{O}_{high}}(x)$ to equal one minus the normalized lowest response between the two highest softmax outputs of $\mathcal{L}$ at $x$, or Margin, as this is a common measure of prediction uncertainty.
2.1 Preliminaries
Notations. Let $P$ denote the underlying probability distribution of data ${\mathbb{X}}$. Assume that a true labeling function $f:{\mathbb{X}}\to{\mathbb{Y}}$ exists. Let ${\mathbb{U}}\subseteq{\mathbb{X}}$ denote the unlabeled set of points, and ${\mathbb{L}}\subseteq{\mathbb{X}}$ the labeled set, such that ${\mathbb{X}}={\mathbb{U}}\cup{\mathbb{L}}$. Here $\left|{\mathbb{L}}\right|=b\leq m$ is the annotation budget where $\left|{\mathbb{X}}\right|=m$.
Let $B_{\delta}\left(x\right)=\left\{x^{\prime}\in{\mathbb{X}}:\left\lVert x^{\prime}-x\right\rVert_{2}\leq\delta\right\}$ denote a ball of radius $\delta$ centered at $x$. Let $C\equiv C({\mathbb{L}},\delta)=\bigcup_{x\in{\mathbb{L}}}B_{\delta}(x)$ denote the region covered by $\delta$-balls centered at the labeled examples in ${\mathbb{L}}$. We call $C({\mathbb{L}},\delta)$ the covered region, where $P\left(C\right)$ denotes its probability. Let $f^{N}$ denote the 1-NN classifier, and $\mathcal{L}$ denote our current learner – a 1-NN classifier based on ${\mathbb{L}}$.
Definition 2.1.
We say that a ball $B_{\delta}(x)$ is pure if $\forall x^{\prime}\in B_{\delta}(x):f(x^{\prime})=f(x)$.
Definition 2.2.
We define the purity of $\delta$ as
$\pi\left(\delta\right)=P\left(\left\{x\in{\mathbb{X}}:B_{\delta}\left(x\right)\text{ is pure}\right\}\right).$
Notice that $\pi(\delta)$ is monotonically decreasing, as can be readily verified.
In [39] it is shown that the generalization error of the 1-NN classifier $f^{N}$ is bounded as follows
${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\left(1-P(C({\mathbb{L}},\delta))\right)+(1-\pi(\delta)).$ (2)
Subsequently, an algorithm is proposed that minimizes the first term in (2) by maximizing the coverage probability $P(C({\mathbb{L}},\delta)$, while ignoring the second term that is assumed to be fixed.
Below, we begin by refining the bound, focusing more closely on its second term $(1-\pi(\delta))$. This is used in Section 3 to device an improved algorithm that minimizes simultaneously both terms of the refined error bound.
2.2 Refined error bound
Define the indicator random variable $I_{\delta}(x)=\mathbbm{1}_{\{x\in{\mathbb{X}}:B_{\delta}(x)\text{ is pure}\}}$. By definition,
$\pi\left(\delta\right)=P\left(\left\{x\in{\mathbb{X}}:B_{\delta}\left(x\right)\text{ is pure}\right\}\right)={\mathbb{E}}[I_{\delta}(x)].$
Since the distribution of ${\mathbb{X}}$ is not known apriori, we use the empirical distribution to approximate the expected value of this random variable. In accordance, given a labeled set ${\mathbb{L}}=\left\{x_{i}\right\}_{i=1}^{b}$:
$\begin{split}\hat{\pi}(\delta)&=\hat{{\mathbb{E}}}[\mathbbm{1}_{\{x\in{\mathbb{L}}:B_{\delta}(x)\text{ is pure}\}}]=\frac{1}{b}{\mathbb{E}}[\sum_{i=1}^{b}\mathbbm{1}_{\{B_{\delta}(x_{i})\text{ is pure}\}}]=\frac{1}{b}\sum_{i=1}^{b}{\mathbb{E}}[\mathbbm{1}_{\{B_{\delta}(x_{i})\text{ is pure}\}}]\\ &=\frac{1}{b}\sum_{i=1}^{b}P(B_{\delta}(x_{i})\text{ is pure}).\end{split}$ (3)
With this approximation
${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big{[}1-P(C({\mathbb{L}},\delta))\Big{]}+\Big{[}1-\frac{1}{b}\sum_{i=1}^{b}P(B_{\delta}(x_{i})\text{ is pure})\Big{]}+\varepsilon.$ (4)
where $\varepsilon$ bounds the error of the empirical approximation in (3).
Note that the refined bound in (4) depends on the purity separately at each labeled point in ${\mathbb{L}}$. We further refine this bound by allowing the fixed raidus $\delta$ to be chosen individually, where $\delta_{i}$ denotes the radius of point $x_{i}\in{\mathbb{L}}$. Let $\Delta=[\delta_{i}]_{i=1}^{b}$ denote the list of individual radii corresponding to ${\mathbb{L}}=\{x_{i}\}_{i=1}^{b}$. The cover defined by $\Delta$ is $C({\mathbb{L}},\Delta)=\bigcup_{(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta}B_{\delta_{i}}(x_{i})$.
We consider the normalized Nearest Neighbor (nNN) classification algorithm, where the nearest neighbor is chosen by solving $f^{N}(x))=\text{argmin}_{x_{i}\in{\mathbb{L}}}\frac{d(x,x_{i})}{\delta_{i}}$. The bound on the error in (4) becomes
${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big{[}1-P(C({\mathbb{L}},\boldsymbol{\Delta}))\Big{]}+\Big{[}1-\frac{1}{b}\sum_{i=1}^{b}P(B_{\delta_{i}}(x_{i})\text{ is pure})\Big{]}+\varepsilon.$ (5)
2.3 Competence score $\boldsymbol{{S_{\mathcal{L}}}}$
In order to obtain a useful competence score [^1] When the labeled set ${\mathbb{L}}$ is large, it may be possible to set aside a validation set to directly estimate the competence of learner $\mathcal{L}$, but this is not feasible when the labeled set ${\mathbb{L}}$ is small., which can be effectively used in objective function (1), we require that it satisfies the following conditions:
  1. (i)
    Lie in the range $[0,1]$.
  2. (ii)
    Use all the data.
  3. (iii)
    Monotonically increase with the competence of learner $\mathcal{L}$.
In accordance, we propose to use the probability of coverage:
${S_{\mathcal{L}}}({\mathbb{L}},\Delta)=\frac{1+e^{-k(1-a)}}{1+e^{-k(P(C({\mathbb{L}},\Delta))-a)}}$ (6)
This function follows a logistic growth curve, with parameters defined as follows: $a\in(0,1)$ is the midpoint and $k\in(0,\infty)$ determines the curve’s steepness.
First, we note that this score satisfies the conditions stated above for the nNN classifier: (i)-(ii), as it is a function in $[0,1]$ that depends on all the data ${\mathbb{X}}$, while (iii) follows from the error bound (5). Below, we prove that it is also monotonically increasing with budget $b$.
More specifically, prior work showed empirically and theoretically that the training budget $b=|{\mathbb{L}}|$ provides a useful clue for the desirable trade-off between ${{\mathcal{O}_{low}}}(x)$ and ${{\mathcal{O}_{high}}}(x)$ in (1), which is captured by score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ defined in (6). We next show that ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is monotonically increasing with $b$. This implies that definition (6) agrees with empirical evidence in prior work.
Proposition.
For two labeled sets ${\mathbb{L}},{\mathbb{L}}^{\prime}$, if ${\mathbb{L}}\subseteq{\mathbb{L}}^{\prime}$ then ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)\leq{S_{\mathcal{L}}}({\mathbb{L}}^{\prime},\Delta^{\prime})$.
Proof.
First, we note that
${\mathbb{L}}\subseteq{\mathbb{L}}^{\prime}\implies C({\mathbb{L}},\Delta)=\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\subseteq\left(\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)\cup\left(\bigcup_{x_{i}\in{\mathbb{L}}^{\prime}\setminus{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)=C({\mathbb{L}}^{\prime},\Delta^{\prime})$
Since probability is monotonic, it follows that $P(C({\mathbb{L}},\Delta))\leq P(C({\mathbb{L}}^{\prime},\Delta^{\prime}))$. From (6), and by the definition of the logistic function using the provided parameters limits, ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is monotonically increasing in ${\mathbb{L}}$, which implies that ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)\leq{S_{\mathcal{L}}}({\mathbb{L}}^{\prime},\Delta^{\prime})$. ∎
3 Our method: DCoM
The error bound in (5) guides our method for active query selection, which aims to minimize this bound by maximizing the coverage $P(C({\mathbb{L}},\Delta))$ and individual purity $\pi(\delta_{i})=P(B_{\delta_{i}}(x_{i})\text{ is pure})$. Recall that ${\mathbb{L}}$ denotes the set of labeled points, $\delta_{i}$ the influence radius of $x_{i}\in{\mathbb{L}}$, $\Delta=[\delta_{i}]_{i=1}^{b}$, and
$C({\mathbb{L}},\Delta)=\bigcup_{(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta}B_{\delta_{i}}(x_{i}),\qquad\qquad\pi(\delta_{i})=P\left(\left\{x\in B_{\delta_{i}}\left(x_{i}\right):f(x)=f(x_{i})\right\}\right)$
Since it was shown in [39] that maximizing $P(C({\mathbb{L}},\delta))$ is already NP-hard, we opt for developing a greedy algorithm for actively selecting the query set, to be termed DCoM.
3.1 Active learning method
The algorithm involves several steps:
Initialization: Define a suitable embedding space for the data and initialize essential parameters.
If ${\mathbb{L}}\neq\emptyset$, train model ${\mathbb{M}}$ on ${\mathbb{L}}$, and compute $\boldsymbol{\delta}$-expansion on the elements of ${\mathbb{L}}$. Otherwise, $\Delta=[\delta_{i}=\delta_{0}]_{i=1}^{b}$.
Repeat /* iterative active learning steps */
  1. (i)
    Active query selection: determine which examples to select for annotation.
  2. (ii)
    Model training: obtain labels and train a deep learning model ${\mathbb{M}}$.
  3. (iii)
    $\boldsymbol{\delta}$-expansion: use ${\mathbb{M}}$ to update $\delta_{i}$ for the recently selected examples.
Until the accumulated budget is exhausted.
We next provide a detailed description of each step.
Initialization: We begin by seeking a suitable embedding space using ${\mathbb{U}}\cup{\mathbb{L}}$, where distances are intended to be inversely correlated with semantic similarity. Self-supervised or representation learning may be used to this end. We choose an initial radius $\delta_{0}$ as described in [39].
(i) Active query selection: In this phase (see Alg 1 below for pseudo-code), the algorithm selects $q$ examples from ${\mathbb{U}}$ to send for labeling, where $q$ denotes the query set size. To this end, it first computes several parameters:
Algorithm 1 DCoM, Active sampling
1:  Input: unlabeled pool ${\mathbb{U}}$, labeled pool ${\mathbb{L}}$, query size $q$, list of ball sizes $\Delta$, trained model ${\mathbb{M}}$ if ${\mathbb{L}}\neq\emptyset$
2:  Output: a set of points to query, and the coverage of ${\mathbb{L}}$
3:  $\forall x\in{\mathbb{U}}$, $M(x)\leftarrow$ normalized margin between the two highest softmax outputs of ${\mathbb{M}}$ if ${\mathbb{L}}\neq\emptyset$, 1 otherwise
4:  Compute $P(C({\mathbb{L}},\Delta))$
5:  Compute ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$
6:  $\delta_{\text{avg}}\leftarrow$ Average($\Delta$)
7:  $G=(V={\mathbb{U}}\cup{\mathbb{L}},E=\{(x,x^{\prime}):x^{\prime}\in B_{\delta_{\text{avg}}}(x)\})$
8:  for $(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta$ do
9:     $\forall x\in B_{\delta_{i}}(x_{i})$: $\forall e=(v,x)\in E$, remove $e$eqparboxCOMMENT# DON’T COVER POINTS TWICE
10:     $\forall e=(x_{i},v)\in E$, remove $e$
11:  end for
12:  ${{\textit{Q}}}\leftarrow\emptyset$
13:  for $i=1,...,q$ do
14:     $\forall x\in{\mathbb{U}}$, $\text{ODR}(x)\leftarrow$Out-Degree Rank
15:     $\forall x\in{\mathbb{U}}$, $R(x)\leftarrow{S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))$ + (1 - ${S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x)$
16:     $x_{max}\leftarrow\{\text{argmax}_{x\in{\mathbb{U}}}R(x)\}$
17:     ${\textit{Q}}\leftarrow{{\textit{Q}}}\cup\{x_{max}\}$
18:     $\forall x\in B_{\delta_{avg}}(x_{max})$: $\forall e=(v,x)\in E$, remove $e$
19:     $M(x_{max})\leftarrow\text{1}$
20:  end for
21:  return  Q, $P(C({\mathbb{L}},\Delta))$
  • $M(x),~{}\forall x\in{\mathbb{U}}$: the normalized margin between the two highest softmax outputs using the last trained model if ${\mathbb{L}}\neq\emptyset$, 1 otherwise.
  • $\Delta$: $\Delta=[\delta_{i}=\delta_{0}]_{i=1}^{b}$ if ${\mathbb{L}}=\emptyset$, otherwise use $\Delta$ from the last AL iteration.
  • $\delta_{\text{avg}}$: the average of $\Delta$.
  • $P(C({\mathbb{L}},\Delta))$: the probability of coverage of the current ${\mathbb{L}}$ with its corresponding $\Delta$.
  • ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$: the competence score of objective function (6) if ${\mathbb{L}}\neq\emptyset$, 0 otherwise.
  • $\{G_{\delta_{\text{avg}}}=(V,E)\}$: a directed adjacency graph where nodes represent samples and edges connect pairs of nodes if the distance between them in the embedding space is smaller than $\delta_{\text{avg}}$. In order not to cover points more than once, $\forall x_{i}\in{\mathbb{L}}$, $\forall x\in B_{\delta_{i}}(x_{i})$ and $\forall{v}\in{V}$ – prune incoming edges $e=(v,x)\in E$. Additionally, we prune all outgoing edges $e=(x_{i},v)\in E$.
Next, $q$ points are to be selected from ${\mathbb{U}}$ in a greedy manner as follows:
  1. 1.
    $\forall x\in{\mathbb{U}}$, compute the Out-Degree-Rank $\text{ODR}(x)$.
  2. 2.
    $\forall x\in{\mathbb{U}}$, compute a ranking score $R(x)={S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))+(1-{S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x)$.
  3. 3.
    Select the vertex $x_{\text{max}}$ with the highest score, $x_{\text{max}}=\text{argmax}_{x\in{\mathbb{U}}}R(x)$.
  4. 4.
    Remove all incoming edges to $x_{\text{max}}$ and its neighbors, implying that $ODR(x_{\text{max}})=0$.
  5. 5.
    Set $M(x_{\text{max}})=1$.
(ii) Model training:
  1. 1.
    Obtain labels for query set Q, returned in step (i) of Active sampling.
  2. 2.
    Remove set Q from ${\mathbb{U}}$ and add it to ${\mathbb{L}}$.
  3. 3.
    Train the given model ${\mathbb{M}}$ using ${\mathbb{L}}$ (and ${\mathbb{U}}$ if the learner is semi-supervised).
(iii) $\boldsymbol{\delta}$-expansion for new samples. In this step (see Alg 2 below for pseudo-code), we compute a suitable $\delta_{i}$ for each new labeled example $x_{i}\in{\textit{Q}}$, and add it to $\Delta$. First, we use the updated model $M$ to predict labels for all the points in the unlabeled set ${\mathbb{U}}$. We also compute a purity threshold $\tau=m\cdot P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))+b$, where $m$ and $b$ are hyperparameters and $P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))$ the probability of coverage using the old labeled set - has been computed in the first Active sampling step. For each new labeled example $v\in{\textit{Q}}$, we search for the largest radius $\delta_{\text{opt}}$ such that its purity $P\left(\left\{x\in B_{\delta_{\text{opt}}}\left(v\right):f(x)=f(v)\right\}\right)$ is still larger than $\tau$. Assuming that $\pi(B_{\delta}(x))$ a monotonic with $\delta$, we can leverage binary search.
Algorithm 2 DCoM, $\boldsymbol{\delta}$-expansion
1:  Input: unlabeled pool ${\mathbb{U}}$, labeled pool ${\mathbb{L}}$, query pool Q, maximal ball size $\delta_{\text{max}}$, list of ball sizes $\Delta$ for ${\mathbb{L}}\setminus{{\textit{Q}}}$, trained model ${\mathbb{M}}$ from the current iteration and $P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))$ from Active sampling
2:  Output: updated list of ball sizes $\Delta$
3:  $\tau\leftarrow m\cdot P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))+b$
4:  $\hat{{\mathbb{Y}}}\leftarrow$ The predicted label for each $x\in{\mathbb{U}}$ using the model ${\mathbb{M}}$
5:  for $v\in{{\textit{Q}}}$ do
6:     $\delta_{\text{opt}}\leftarrow\text{argmax}_{\delta}[P\left(\left\{x\in B_{\delta}\left(v\right):f(x)=f(v)\right\}\right)>\tau$]
7:     $\Delta\leftarrow\Delta+[\delta_{\text{opt}}]$
8:  end for
9:  return  $\Delta$
Using a sparse representation of the adjacency graph enables DCoM to handle large datasets efficiently without exhaustive space requirements. The algorithm’s complexity, including adjacency graph construction and sample selection, is discussed in App. C.
4 Empirical evaluation
4.1 Methodology
This section investigates two deep AL frameworks: a fully supervised approach, wherein a deep model is trained exclusively on the annotated dataset ${\mathbb{L}}$ as a standard supervised task, and a semi-supervised framework trained on both the annotated and unlabeled datasets, ${\mathbb{L}}$ and ${\mathbb{U}}$. Our experimental setup is based on the codebase of [24], ensuring a fair comparison among the various AL strategies by using the same network architectures and sharing all relevant experimental conditions.
More specifically, in the following experiments, we trained ResNet-18 [16] on STL-10 [6], SVHN [25], CIFAR-100 [22] and subsets of ImageNet [9] including ImageNet-50, ImageNet-100, and ImageNet-200 as they appear in [33]. The hyper-parameters we used are detailed in App. A. In the semi-supervised framework, we used FlexMatch [42] and the code provided by [35]. For the experiments, we followed the parameters specified by [42]. The exact parameters are detailed in App. A. While the ResNet-18 architecture may no longer achieve state-of-the-art results on the datasets examined, it serves as a suitable platform to evaluate the efficacy of AL strategies in a competitive and fair environment, where they have demonstrated benefits.
In our comparisons, we used several active learning strategies (or optimization criteria) for selecting a query set: (1) Randomsampling, (2) Min margin— lowest margin between the two highest softmax outputs, (3) Max entropy— highest entropy of softmax outputs [30], (4) Uncertainty— 1 minus maximum softmax output, (5) DBAL[11], (6) CoreSet[28], (7) BALD[21], (8) BADGE[1], (9) ProbCover[39], and (10) LDM-s[4]. When available, we used the code provided in [24] for each strategy. When it was unavailable, we used the code from the repositories of the corresponding papers. Since [4] doesn’t provide code, we compared our results with the corresponding available results reported in their paper, despite differences in the running setup (see App. D.6).
DCoM  and ProbCover require a suitable data embedding. For STL-10 , SVHN, CIFAR-10 and CIFAR-100 we employed the SimCLR [3] embedding. For the ImageNet subsets we used the DINOv2 [26] embedding. Section 4.3 provides an extensive ablation study, which includes alternative embedding spaces. In all experiments, identical settings and hyper-parameters are used for our method, as detailed in App. A. The robustness of our method to these choices is discussed in App. B.
4.2 Main results
(a) STL-10
(b) SVHN
(c) CIFAR-100
(d) ImageNet-50
(e) ImageNet-100
(f) ImageNet-200
Figure 1: Mean accuracy difference between several AL algorithms and random queries. Positive mean difference indicates that the corresponding AL method is beneficial. The errors bars correspond to standard error, based on 5-10 repetitions (3 for the ImageNet subsets). While some methods perform well only in specific budget regimes, DCoM consistently achieves the best results across all budgets.
Fully supervised framework. We evaluate various AL methods by training a ResNet-18 model from scratch for each different AL strategy, using the labels of the queries selected by the respective method. In each AL iteration, a new model is trained from scratch using the updated labeled set ${\mathbb{L}}$. This process is repeated for several AL iterations. Fig. 1 shows the difference between the accuracy obtained by each AL method, and the accuracy obtained when training a similar network with a random query set of the same size (the baseline of no active learning), for each AL iteration. The final accuracy of each method is shown in Tables 1-2 for the ImageNet-100 and CIFAR-100 datasets, as mean $\pm$ Standard Error (STE). Additional datasets are shown in App. D.
Table 1: Model final accuracy when varying the size of the labeled set ${\mathbb{L}}$ (row) and the respective AL strategy (column), using the ImageNet-100 dataset
$|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin DCoM
200 4.58±0.13 6.72±0.07 4.67±0.16 1.99±0.22 4.58±0.13 2.01±0.06 1.99±0.10 1.83±0.06 1.84±0.12 6.87±0.18
400 6.61±0.17 9.30±0.23 6.42±0.23 4.39±0.39 5.53±0.16 3.50±0.32 4.17±0.33 3.65±0.25 4.71±0.10 10.59±0.26
600 8.11±0.42 12.71±0.42 8.01±0.20 6.42±0.24 6.45±0.37 4.93±0.29 5.51±0.33 4.96±0.22 6.46±0.12 13.61±0.15
1,000 11.43±0.17 17.71±0.27 11.56±0.38 9.42±0.19 9.04±0.66 7.97±0.82 8.39±0.18 7.45±0.18 10.43±0.11 17.53±0.23
3,000 27.59±0.76 34.01±0.15 27.58±0.72 24.96±0.23 19.41±0.32 23.10±0.60 20.97±0.33 20.88±0.52 26.51±0.79 34.85±0.23
5,000 39.25±0.26 41.91±0.20 39.13±0.06 36.50±0.25 29.35±0.39 33.78±0.46 31.89±0.24 31.25±0.21 38.37±0.10 44.09±0.83
7,000 46.88±0.25 48.41±0.42 46.42±0.41 43.49±0.42 38.45±0.12 42.09±0.54 41.19±0.17 40.47±0.83 45.69±0.26 49.22±0.29
11,000 55.31±0.26 54.01±0.15 56.61±0.11 55.29±0.32 50.63±0.33 53.25±0.31 52.52±0.26 53.09±0.29 56.86±0.10 58.45±0.15
15,000 61.69±0.09 57.04±0.43 63.05±0.42 62.19±0.21 58.71±0.05 60.07±0.07 59.76±0.22 59.64±0.30 62.71±0.27 64.57±0.22
19,000 66.39±0.26 59.69±0.16 66.82±0.33 66.19±0.19 64.77±0.32 65.59±0.27 64.65±0.23 64.93±0.23 66.89±0.28 67.60±0.08
23,000 68.92±0.15 61.99±0.41 70.54±0.21 69.62±0.32 68.49±0.06 69.03±0.20 68.47±0.50 67.65±0.14 70.81±0.35 70.21±0.32
Table 2: Same as Table 1, using the CIFAR-100 dataset.
$|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin LDM-s DCoM
200 6.39±0.19 10.79±0.08 6.31±0.09 3.30±0.11 6.39±0.19 2.59±0.05 2.74±0.17 2.61±0.18 5.09±0.37 - 10.59±0.19
400 8.74±0.07 13.53±0.22 8.27±0.14 6.63±0.23 8.62±0.17 4.59±0.23 4.68±0.23 4.41±0.30 7.58±0.17 - 14.11±0.11
600 10.73±0.13 15.86±0.16 10.29±0.27 8.33±0.31 10.39±0.25 6.52±0.38 6.32±0.27 5.94±0.28 10.20±0.16 - 16.66±0.11
1,000 13.49±0.18 19.49±0.20 13.66±0.26 11.42±0.26 12.94±0.21 9.98±0.38 9.31±0.22 8.47±0.35 13.28±0.19 - 20.21±0.07
3,000 24.02±0.27 29.31±0.42 24.24±0.28 22.10±0.11 23.00±0.28 21.93±0.26 20.15±0.33 20.29±0.23 23.71±0.31 - 30.36±0.08
5,000 31.52±0.39 34.64±0.15 32.18±0.33 29.58±0.38 31.41±0.26 30.60±0.40 29.15±0.46 29.60±0.28 31.66±0.23 - 36.95±0.10
7,000 37.69±0.16 38.36±0.13 38.09±0.10 34.36±0.20 38.24±0.16 35.87±0.19 35.11±0.11 35.67±0.35 38.16±0.17 31.85±0.00 41.52±0.25
11,000 45.66±0.22 44.69±0.20 47.00±0.16 44.37±0.28 46.53±0.18 45.23±0.20 45.01±0.17 44.31±0.24 46.37±0.36 40.88±0.01 46.86±0.28
15,000 50.57±0.09 49.58±0.46 52.42±0.17 51.02±0.09 52.19±0.29 51.82±0.16 51.43±0.21 51.30±0.35 52.37±0.19 46.59±0.01 52.33±0.22
19,000 54.95±0.09 52.90±0.14 56.32±0.07 55.55±0.27 56.72±0.31 56.55±0.21 55.76±0.31 55.75±0.14 56.73±0.24 49.41±0.01 56.53±0.07
23,000 57.74±0.14 56.29±0.21 59.60±0.16 59.96±0.35 59.59±0.22 59.79±0.23 59.91±0.08 59.72±0.21 59.88±0.21 52.48±0.00 59.60±0.33
Semi-supervised framework. We evaluate the effectiveness of 5 representative AL strategies with FlexMatch, a competitive semi-supervised learning method. We also evaluate our method, and the results of random sampling (no AL) as baseline. Fig. 2 shows the mean accuracy and STE based on 3 repetitions of FlexMatch, employing labeled sets generated by the different AL strategies. Note that in these experiments, since the query is selected from an unlabeled set, the resulting labeled set ${\mathbb{L}}$ may be unbalanced between the classes, which is especially harmful when the budget is small. This affects all the methods. As a result, the accuracy with random sampling may differ from reported results. Notably, DCoM demonstrates significant superiority over random sampling.
$|{\mathbb{L}}|=400$
$|{\mathbb{L}}|=1,000$
Figure 2: Comparison of AL strategies in a semi-supervised learning framework. Each bar represents the mean test accuracy over 3 repetitions of FlexMatch, trained using a total of 400 (left) and 1,000 (right) labeled examples on the CIFAR-100 dataset (an average of 4 and 10 labeled examples per class respectively). Three rounds of AL are considered, selecting 100 examples in the first round, 300 in the second, and 600 in the third.
4.3 Ablation study
We present a series of experiments where we investigate the choices made by our method, including different feature spaces and different uncertainty-based measures. The results demonstrate the robustness of our method to these choices. Additional choices are investigate in App. D, including the contribution of a dynamic $\delta$ and initialization $\delta_{0}$ (App. D.1). In App. D.3, we evaluate the effectiveness of ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ from (6) in capturing competence.
Different feature spaces. As discussed in Section 3.1, our approach relies on the existence of a suitable embedding space, where distance is indicative of semantic dissimilarity. We now repeat the basic fully-supervised experiments while varying the embedding employed by DCoM, considering various popular self-supervised representations. Results are reported in Fig. 4, showing that DCoM consistently delivers the best performance regardless of the specific embedding used.
Breaking down the algorithm into its components. We evaluate the contribution of each component in objective function (1) separately. Results are shown in Fig. 4, wherein DPC denotes the Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM, and Margin denotes DCoM’s choice for ${{\mathcal{O}_{high}}}(x)$. Clearly, the weighted objective function of DCoM yields superior results. Similar behavior is observed for several other datasets and other selections of ${{\mathcal{O}_{high}}}(x)$, see App. D.4.
(a) MOCOv2+
(b) Barlow Twins
Figure 3: Performance over CIFAR-100, using two different embedding spaces for DCoM and ProbCover, see details in the caption of Fig. 1. Clearly, DCoM consistently achieves the best results. Results using another embedding space are provided in App. D.5.
(c) CIFAR-10
Figure 4: Performance evaluation (mean accuracy and STE) when optimizing each component of objective function (1) separately and together.
Different uncertainty-based methods. We now evaluate our choice for ${{{\mathcal{O}_{high}}}(x)}$ in objective function (1), investigating the alternative methods of min margin, entropy of the min max activation, and uncertainty (see details in Section 4.1). Additionally, we use a score reminiscent of the AL method BADGE [11] and termed Gradient norm, which computes the norm of the gradient between the learner’s last two layers. Once again, the results of DCoM show robustness to this choice, with insignificant differences between the different uncertainty scores, see Fig. 5. Consequently, we prioritize computational efficiency and opt for the fastest-to-compute score, which is min-margin.
Table 3: Model accuracy for different sizes of labeled set ${\mathbb{L}}$ and several uncertainty-based methods for ${\mathbb{F}}(x)$
$|{\mathbb{L}}|$ Random Entropy Margin Uncertainty Gradient norm
200 6.39±0.19 10.71±0.30 10.59±0.19 10.44±0.26 10.33±0.21
400 8.74±0.07 14.35±0.24 14.11±0.11 14.72±0.20 14.59±0.19
600 10.73±0.13 17.07±0.19 16.66±0.11 16.75±0.34 16.83±0.17
1,000 13.49±0.18 19.98±0.22 20.21±0.07 20.40±0.21 20.03±0.17
3,000 24.02±0.27 29.89±0.11 30.36±0.08 30.30±0.11 30.62±0.52
5,000 31.52±0.39 36.08±0.13 36.95±0.10 36.06±0.18 36.12±0.46
7,000 37.69±0.16 41.18±0.25 41.52±0.25 41.80±0.08 41.23±0.65
11,000 45.66±0.22 47.23±0.09 46.86±0.28 46.99±0.27 46.84±0.28
15,000 50.57±0.09 52.67±0.15 52.33±0.22 52.43±0.14 52.21±0.22
19,000 54.95±0.09 56.64±0.27 56.53±0.07 56.47±0.06 56.35±0.28
23,000 57.74±0.14 59.50±0.18 59.60±0.33 59.68±0.24 59.45±0.06
Figure 5: Similar to Fig. 1 and Tables 1-2, using only DCoM while leveraging 4 different variants for its underlying uncertainty score (see text). The performance differences between the variants are small.
5 Summary
We investigate the challenge of Active Learning (AL) in a multi-budget setting. Our approach is motivated by a new bound on the generalization error of the Nearest Neighbor classifier, and aims to minimize this bound. Additionally, the sampling strategy is dynamically adjusted during training using a new competence score, also motivated by the same bound. We validate our method empirically in both supervised and semi-supervised frameworks across various datasets. The results show that DCoM significantly outperforms alternative methods across all budgetary constraints.
Appendix
Appendix A Implementation details
The source code used in this study is available at https://github.com/avihu111/TypiClust.
The code is based on the GitHub repositories of Yehuda et al. [39] and Munjal et al. [24].
DCoM necessitates the computation of the initial delta for each dataset representation. The values utilized in our experiments can be found in Table 4. The $\delta$ values of CIFAR-10 and CIFAR-100 are sourced from [39]. Information on the computation of $\delta$ values for other datasets can be located in App. D.1.
Table 4: Initial delta values used in experiments
Dataset SSL method Initial $\delta$
STL-10 SimCLR 0.55
SVHN SimCLR 0.4
CIFAR-10 SimCLR 0.75
CIFAR-100 SimCLR 0.65
ImageNet-50 DINOv2 0.75
ImageNet-100 DINOv2 0.7
ImageNet-200 DINOv2 0.65
CIFAR-100 MOCOv2+ 0.5
CIFAR-100 BYOL 0.5
CIFAR-100 Barlow Twins 0.55
Additionally, throughout our experiments we used identical algorithm parameters, distinguishing between those for datasets with smaller class amounts (10) and larger ones (50+). These parameters encompass the adaptive purity threshold $\tau$ from Alg 2 and the logistic function parameters for the competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ in (6), $(a,k)$. Specifically, we employed $\tau=0.2~{}\text{cover}+0.4$, with logistic parameters $a=0.9$ and $k=30$ for CIFAR-10, SVHN, and STL-10, and $a=0.8$ and $k=30$ for CIFAR-100 and ImageNet subsets. The $\delta$ resolution for the binary search in DCoM was set to $0.05$. Our ablation study, detailed in App. B, demonstrates that these selections have negligible impact on performance.
A.1 Supervised training
When training on STL-10, SVHN, CIFAR-10 and CIFAR-100 datasets, we utilized a ResNet-18 architecture trained for 200 epochs. Our optimization strategy involved using an SGD optimizer with a Nesterov momentum of 0.9, weight decay set to 0.0003, and cosine learning rate scheduling starting at a base rate of 0.025. Training was performed with a batch size of 100 examples, and we applied random cropping and horizontal flips for data augmentation. For an illustration of these parameters in use, refer to [24].
When training ImageNet-50, we used the same hyper-parameters, only changing the base learning rate to $0.01$, the batch size to $50$ and the epoch amount to $50$.
When training ImageNet-100/200, we used the same hyper-parameters as ImageNet-50, only changing the epoch amount to $100$.
The train and test partitions followed the original sets from the corresponding paper, with $10\%$ of the train set used for validation in all datasets except for ImageNet-200, where $5\%$ was used for validation.
A.2 Semi-supervised training
When training FlexMatch [42], we used the semi supervised framework by [35], and the AL framework by [24]. All experiments involved 3 repetitions.
We relied on the standard hyper-parameters used by FlexMatch [42]. Specifically, we trained WideResNet-28 for 512 epochs using the SGD optimizer, with $0.03$ learning rate, $64$ batch size, $0.9$ SGD momentum, $0.999$ EMA momentum, $0.001$ weight decay and $2$ widen factor.
A.3 Self-supervised feature extraction
STL-10, SVHN, CIFAR-10, CIFAR-100. To extract semantically meaningful features, we trained SimCLR using the code provided by [33] for STL-10, SVHN, CIFAR-10 and CIFAR-100. Specifically, we used ResNet-18 [16] with an MLP projection layer to a $128$-dim vector, trained for 512 epochs. All the training hyper-parameters were identical to those used by SCAN (all details can be found in [33]). After training, we used the $512$ dimensional penultimate layer as the representation space.
Representation learning: ImageNet-50/100/200. We extracted features from the official (ViT-S/14) DINOv2 weights pre-trained on ImageNet [26]. We employed the L2-normalized penultimate layer for the embedding, which has a dimensionality of $384$.
Ablation embeddings – CIFAR-100. We extracted more features as BYOL [13], MOCOv2+ [17], and Barlow Twins [41] using pre-trained weights from [8].
Appendix B Hyper-parameters exploring
As described before, DCoM has 3 hyper-parameters: $a,k\text{and }\tau$. When running the main experiments described in Section 4.2, we maintained consistency by employing identical parameters for STL-10, SVHN, and CIFAR-10, as well as for CIFAR-100 and the subsets of ImageNet. This approach helps to demonstrate that the choice of the hyper-parameters is not critical and does not significantly affect the results. Here, we conduct experiments with DCoM using STL-10 and CIFAR-100 datasets and several values for $a$ and $k$ from the definition of the competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ in (6). We observe that while certain choices may be slightly better than others, all choices yield results that are very close to each other, and DCoM consistently outperforms or matches the performance of previous methods.
(a) STL-10
(b) CIFAR-100
Figure 6: Evaluation of DCoM across different settings of the logistic function hyper-parameters a and k. Each plot displays the mean and standard error over $3-5$ repetitions. The findings indicate that fine-tuning these parameters has negligible effects on performance.
Appendix C Time and space complexity of DCoM
For the complexity analysis below, let $n$ represent the number of examples in the combined unlabeled and labeled pool $|{\mathbb{L}}\cup{\mathbb{U}}|$, $d$ the dimension of the data embedding space, $b$ the given budget ($|{\mathbb{L}}|$) and $q$ the size of the query set. As previously stated, DCoM can be divided into three distinct steps:
C.1 Selection of initial $\boldsymbol{\delta}$
Time Complexity: The process begins by creating $t$ adjacency graphs, each corresponding to a different $\delta$ value being evaluated. Here, $t$ denotes the number of $\delta$’s being examined. The time complexity of each graph generation is $O(n^{2}\cdot d)$ time. The computation of purity requires the prediction of pseudo-labels by k-means at a cost of $O(n^{2})$. The complexity of computing ball purity for each example is $O(n^{2})$, so totally each the ball purity computing takes $O(n^{3})$. This adds up to overall time complexity of $O(tdn^{3})$. Although it may seem significant, this step doesn’t occur during the active learning algorithm process. Additionally, it’s possible to confine the data employed for purity computation to the densest points, effectively lowering the complexity to $O(tdn^{2})$. This approach mandates the use of at least the number of points equivalent to the number of classes for purity computation. This technique yields similar $\delta_{0}$ values (consistent for CIFAR-10, CIFAR-100, and ImageNet-50, and with a variation of 0.05 for ImageNet-100/200 and STL-10 and 0.15 for SVHN).
Space complexity: Naively, the space complexity is $O(n^{2})$, which might be impractical for large datasets like ImageNet. However, [39] demonstrates that utilizing a sparse matrix in coordinate list (COO) format and using limited $\delta$ values, the space complexity becomes $O(|E|)$, where $E$ represents the set of edges in the graph. Although $O(|E|)$ remains $O(n^{2})$ in the worst-case scenario, in practice, the average vertex degree with a limited radius is less than $n$.
C.2 Customizing $\boldsymbol{\delta}$ values
Time complexity: Each iteration of DCoM starts with customizing $\delta$ values for each labeled example from the last active step. In the worst case, there are $|{\mathbb{L}}|$ examples. Customizing $\delta$ involves building the adjacency graph using $\delta_{\text{max}}$, running all examples on the model to obtain their predictions, and applying binary search over $\delta$ values between 0 and $\delta_{\text{max}}$ for each example from the last active step. The binary search is performed on a continuous parameter ($\delta$) with a search resolution $r$. This implies that there are $\frac{\delta_{\text{max}}}{r}$ available values for $\delta$. Overall, this process takes $O(dn^{2}+n+|{\mathbb{L}}|\cdot\log(\frac{\delta_{\text{max}}}{r}))=O(dn^{2}+b\cdot\log(\frac{\delta_{\text{max}}}{r}))$. In practical terms, due to the vectorization of these processes, it requires approximately 27 minutes to construct an adjacency graph for ImageNet-200 and update deltas for 1000 examples on a single CPU.
Space complexity: Similar to the space complexity analysis in App. C.1, which is $O(n^{2})$.
C.3 Active sampling
Time complexity: Active sampling involves balancing model min-margin and example density based on the current adjacency graph. Initially, the margin of each example is computed using the softmax output from the model’s last layer, requiring $O(n)$ time. Subsequently, the current adjacency is created using $\delta_{\text{avg}}$, which is the average over the ${\mathbb{L}}$ $\delta$ values list - $\Delta$. As discussed in App. C.1, this step has a time complexity of $O(dn^{2})$. Following this, the iterative process for selecting a single sample includes the following steps:
  • Calculating node degrees – $O(|E|)$ time.
  • Finding the node with maximal degree – $O(n)$ time.
  • Removing incoming edges from the graph for covered points – $O(|E|)$ time.
Samples are iteratively selected from the current sparse graph, with incoming edges to newly covered samples being removed. Unlike adjacency graph creation, sample selection cannot be parallelized, as each step depends on the previous one. Overall, the time complexity for selecting a sample is $O(|E|+n)$, resulting in a worst-case overall complexity of $O(dn^{2})$. However, as more points are selected, the removal of edges speeds up the selection of later samples. Practically, it consumes about 15 minutes to construct an adjacency graph and select 1000 samples from ImageNet-200 on a single CPU.
Space complexity: Similar to the previous space complexity analysis – $O(n^{2})$.
Appendix D Additional empirical results
D.1 $\delta_{0}\,\,\,$ Initialization
To set the initial value $\delta_{0}$, we adopted the method outlined in [39], as detailed in Section 3. Fig. 7 displays the purity function across various $\delta$ values, along with the selected $\delta$ values for each dataset. As mentioned previously, the $\delta$ values for CIFAR-10 and CIFAR-100 are derived from [39].
(a) STL-10
(b) SVHN
(c) ImageNet-50
(d) ImageNet-100
(e) ImageNet-200
Figure 7: $\pi(\delta)$, estimated from the unlabeled data and using k-means algorithm for labeling. The dashed line indicates the highest $\delta$, after which purity drops below $\alpha=0.95$.
D.2 $\Delta\,\,\,\,$ distribution throughout the training process
The experiment in Fig. 8 illustrates the distribution of $\Delta$ throughout the active selection and training phases of DCoM. The significant standard deviation of $\Delta$ compared to the minor standard error underscores the significance of employing a dynamic method that assigns a distinct radius to each ball.
(a) CIFAR-10
(b) CIFAR-100
(c) ImageNet-50
Figure 8: The $\Delta$ distribution through DCoM algorithm. Each plot displays the mean and standard error over 5-10 repetitions. The results indicate that as the emphasis shifts towards maximizing the purity of sample balls, rather than solely focusing on maximum coverage, the radii adapt accordingly. This, combined with the improved performance observed with DCoM, highlights the necessity of the dynamic algorithm in enhancing the effectiveness of active sampling.
D.3 Dynamic algorithm coverage
The analysis presented in Fig. 9 demonstrates a consistent behavior of probability coverage across different datasets during the DCoM process. This indicates that using probability coverage as a measure of progress consistently proves to be a wise decision.
(a) CIFAR-10 and STL-10
(b) CIFAR-100 and ImageNet-50
Figure 9: Comparison between $P(C({\mathbb{L}},\Delta))$ values during DCoM sampling process, across several datasets. Each plot displays the mean and standard error over $3-10$ repetitions. These figures demonstrates a consistent behavior of probability coverage.
D.4 Decomposing the algorithm
In Section 4.3, we present an ablation study where we evaluate the contribution of each component in objective function (1) separately on CIFAR-100 dataset. In Fig. 10 you can see the results over CIFAR-100 and ImageNet-50 datasets.
(a) CIFAR-100
(b) ImageNet-50
Figure 10: Performance evaluation (mean accuracy and STE) when optimizing each component of objective function (1) separately and together. DPC denotes the Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM. Margin is DCoM’s choice for ${{\mathcal{O}_{high}}}(x)$.
In Fig. 11 you can see the same ablation but using entropy as ${{\mathcal{O}_{high}}}(x)$ instead of margin.
(a) CIFAR-10
(b) CIFAR-100
(c) ImageNet-50
Figure 11: Assessment of each algorithm’s objective function individually across multiple datasets. DPC stands for Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM. DCEM represents DCoM with entropy as ${{\mathcal{O}_{high}}}(x)$ instead of margin. Each plot displays the mean and standard error over $5-10$ repetitions. These graphs demonstrate how the weighted objective function consistently produces superior results across all labeled set sizes $|{\mathbb{L}}|$.
D.5 Different embedding space
In Section 4.3, we present an ablation study where we repeat the basic fully-supervised experiments while varying the embedding employed by DCoM. In Fig. 12 you can see the same ablation using BYOL [13] representation.
Figure 12: Performance over CIFAR-100, using BYOL feature space over DCoM and ProbCover, see details in the caption of Fig. 1. Clearly, DCoM consistently achieves the best results.
D.6 Comparison per dataset
Tables 5-6 present the empirical accuracy values of several datasets using different active learning algorithms. In each active step, a new learner is trained from scratch over all available labeled data, and the results are presented as the [mean $\pm$ STE] over $5-10$ repetitions (3 for ImageNet subsets). As mentioned earlier, the results for LDM-s algorithm were acquired from the study by [4]. Their lack of code provision suggests discrepancies in the running setup. Nonetheless, the disparity between their method and random selection in their setup is smaller than that observed in ours, implying the superiority of our method. The table includes the results for all datasets, with the name of each dataset below it.
Table 5: Model accuracy for different sizes of labeled set ${\mathbb{L}}$ and AL strategies
$|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin LDM-s DCoM
10 14.53±0.60 20.39±0.23 15.17±0.94 10.81±0.61 14.65±0.74 12.99±0.77 12.78±0.72 11.74±0.69 14.39±0.88 - 19.72±0.40
20 16.82±0.44 24.64±0.73 17.85±0.77 13.79±0.89 16.09±0.74 16.38±0.52 17.07±0.62 17.00±0.49 17.33±0.85 - 22.28±0.75
30 18.82±0.71 27.19±0.25 20.38±0.67 16.25±0.83 18.22±0.93 18.00±0.68 19.26±0.44 18.34±0.50 18.55±0.87 - 26.98±0.35
40 20.89±0.55 29.46±0.38 22.19±0.45 17.93±0.67 19.79±0.61 19.00±0.70 19.98±0.60 20.30±0.35 20.61±0.68 - 28.91±0.46
50 22.75±0.42 29.80±0.21 22.41±0.35 18.36±0.54 21.26±0.63 21.63±0.70 22.17±0.48 21.06±0.45 22.00±0.61 - 30.39±0.12
200 31.53±0.29 38.83±0.18 32.10±0.42 27.95±0.29 30.91±0.76 30.69±0.27 29.51±0.42 28.91±0.50 31.83±0.48 - 38.86±0.40
400 38.64±0.36 45.36±0.51 37.92±0.37 35.10±0.33 37.85±0.70 36.48±0.46 36.71±0.38 36.01±0.32 38.65±0.62 - 45.72±0.27
600 43.66±0.44 49.13±0.36 42.69±0.35 40.75±0.29 42.48±0.39 41.96±0.24 40.75±0.23 41.19±0.53 43.75±0.22 - 49.71±0.39
800 46.84±0.29 51.49±0.37 46.34±0.32 44.12±0.49 46.21±0.25 45.50±0.17 43.89±0.31 44.93±0.42 46.97±0.40 - 52.10±0.58
1,400 54.35±0.27 56.33±0.26 54.06±0.42 53.08±0.34 52.31±0.31 53.13±0.28 52.93±0.34 53.28±0.50 54.97±0.35 50.30±0.00 57.54±0.10
2,600 62.93±0.31 62.16±0.18 64.67±0.31 64.19±0.24 61.94±0.40 64.98±0.59 64.62±0.46 63.53±0.44 64.54±0.45 57.01±0.01 65.76±0.21
3,800 68.61±0.44 66.83±0.25 71.36±0.32 70.77±0.53 68.41±0.34 71.90±0.36 71.30±0.36 70.55±0.25 71.60±0.21 61.30±0.01 71.84±0.36
5,000 73.61±0.23 70.95±0.23 75.93±0.21 75.86±0.27 72.92±0.36 75.82±0.31 76.13±0.31 75.24±0.47 75.46±0.31 64.09±0.00 76.36±0.05
6,200 75.95±0.18 73.51±0.30 78.90±0.27 77.92±0.36 77.00±0.19 78.84±0.27 78.37±0.20 78.54±0.26 78.62±0.22 65.79±0.00 78.66±0.16
7,400 78.49±0.14 75.71±0.35 80.88±0.22 80.94±0.16 79.39±0.30 81.19±0.15 81.02±0.19 81.10±0.24 80.76±0.29 67.28±0.01 81.30±0.34
CIFAR-10 dataset
$|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin LDM-s DCoM
200 6.39±0.19 10.79±0.08 6.31±0.09 3.30±0.11 6.39±0.19 2.59±0.05 2.74±0.17 2.61±0.18 5.09±0.37 - 10.59±0.19
400 8.74±0.07 13.53±0.22 8.27±0.14 6.63±0.23 8.62±0.17 4.59±0.23 4.68±0.23 4.41±0.30 7.58±0.17 - 14.11±0.11
600 10.73±0.13 15.86±0.16 10.29±0.27 8.33±0.31 10.39±0.25 6.52±0.38 6.32±0.27 5.94±0.28 10.20±0.16 - 16.66±0.11
1,000 13.49±0.18 19.49±0.20 13.66±0.26 11.42±0.26 12.94±0.21 9.98±0.38 9.31±0.22 8.47±0.35 13.28±0.19 - 20.21±0.07
3,000 24.02±0.27 29.31±0.42 24.24±0.28 22.10±0.11 23.00±0.28 21.93±0.26 20.15±0.33 20.29±0.23 23.71±0.31 - 30.36±0.08
5,000 31.52±0.39 34.64±0.15 32.18±0.33 29.58±0.38 31.41±0.26 30.60±0.40 29.15±0.46 29.60±0.28 31.66±0.23 - 36.95±0.10
7,000 37.69±0.16 38.36±0.13 38.09±0.10 34.36±0.20 38.24±0.16 35.87±0.19 35.11±0.11 35.67±0.35 38.16±0.17 31.85±0.00 41.52±0.25
11,000 45.66±0.22 44.69±0.20 47.00±0.16 44.37±0.28 46.53±0.18 45.23±0.20 45.01±0.17 44.31±0.24 46.37±0.36 40.88±0.01 46.86±0.28
15,000 50.57±0.09 49.58±0.46 52.42±0.17 51.02±0.09 52.19±0.29 51.82±0.16 51.43±0.21 51.30±0.35 52.37±0.19 46.59±0.01 52.33±0.22
19,000 54.95±0.09 52.90±0.14 56.32±0.07 55.55±0.27 56.72±0.31 56.55±0.21 55.76±0.31 55.75±0.14 56.73±0.24 49.41±0.01 56.53±0.07
23,000 57.74±0.14 56.29±0.21 59.60±0.16 59.96±0.35 59.59±0.22 59.79±0.23 59.91±0.08 59.72±0.21 59.88±0.21 52.48±0.00 59.60±0.33
CIFAR-100 dataset
$|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin DCoM
200 8.25±0.65 12.33±0.34 8.82±0.18 4.46±0.31 8.31±0.52 4.26±0.26 4.11±0.09 4.26±0.24 6.50±0.94 12.27±0.04
400 12.18±0.29 16.64±0.29 12.73±0.42 9.49±0.84 9.40±0.59 7.74±0.41 7.01±0.37 6.45±0.33 11.48±0.72 18.03±0.21
600 14.85±0.53 20.38±0.47 15.69±0.38 11.21±0.41 10.68±0.35 10.38±0.27 11.28±0.54 10.34±0.46 14.74±0.35 21.77±0.62
1,000 19.71±0.53 25.76±0.34 20.91±0.43 14.85±0.35 13.47±0.32 16.30±0.60 16.52±0.12 15.10±0.65 19.47±0.45 26.88±0.00
3,000 39.06±0.21 39.98±0.23 38.51±0.15 32.40±0.40 28.80±0.45 32.52±0.63 32.97±0.34 31.60±0.60 37.00±0.65 43.40±0.29
5,000 48.33±0.27 45.08±0.16 49.47±0.26 45.04±0.47 40.10±0.31 42.05±0.29 41.65±0.22 42.41±0.52 47.24±0.50 50.76±0.08
7,000 55.12±0.09 47.34±0.44 55.25±0.34 53.48±0.38 49.48±0.19 51.30±0.68 49.74±0.21 49.67±0.46 54.78±0.29 55.17±0.34
11,000 63.46±0.15 52.79±0.17 64.68±0.15 63.65±0.32 60.13±0.11 60.10±0.50 60.89±0.34 60.58±0.25 63.90±0.22 63.51±0.31
15,000 69.42±0.20 55.04±0.06 70.27±0.25 70.08±0.21 67.38±0.17 67.49±0.19 66.34±0.16 66.65±0.23 69.36±0.24 70.19±0.45
19,000 72.54±0.21 62.66±0.19 73.70±0.29 73.63±0.21 71.87±0.14 72.78±0.41 71.55±0.27 71.85±0.17 73.52±0.23 74.29±0.50
23,000 75.99±0.16 70.06±0.27 76.65±0.11 77.17±0.06 75.30±0.24 76.37±0.29 75.73±0.25 75.64±0.27 77.12±0.30 76.68±0.20
ImageNet-50 dataset
$|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin DCoM
200 4.58±0.13 6.72±0.07 4.67±0.16 1.99±0.22 4.58±0.13 2.01±0.06 1.99±0.10 1.83±0.06 1.84±0.12 6.87±0.18
400 6.61±0.17 9.30±0.23 6.42±0.23 4.39±0.39 5.53±0.16 3.50±0.32 4.17±0.33 3.65±0.25 4.71±0.10 10.59±0.26
600 8.11±0.42 12.71±0.42 8.01±0.20 6.42±0.24 6.45±0.37 4.93±0.29 5.51±0.33 4.96±0.22 6.46±0.12 13.61±0.15
1,000 11.43±0.17 17.71±0.27 11.56±0.38 9.42±0.19 9.04±0.66 7.97±0.82 8.39±0.18 7.45±0.18 10.43±0.11 17.53±0.23
3,000 27.59±0.76 34.01±0.15 27.58±0.72 24.96±0.23 19.41±0.32 23.10±0.60 20.97±0.33 20.88±0.52 26.51±0.79 34.85±0.23
5,000 39.25±0.26 41.91±0.20 39.13±0.06 36.50±0.25 29.35±0.39 33.78±0.46 31.89±0.24 31.25±0.21 38.37±0.10 44.09±0.83
7,000 46.88±0.25 48.41±0.42 46.42±0.41 43.49±0.42 38.45±0.12 42.09±0.54 41.19±0.17 40.47±0.83 45.69±0.26 49.22±0.29
11,000 55.31±0.26 54.01±0.15 56.61±0.11 55.29±0.32 50.63±0.33 53.25±0.31 52.52±0.26 53.09±0.29 56.86±0.10 58.45±0.15
15,000 61.69±0.09 57.04±0.43 63.05±0.42 62.19±0.21 58.71±0.05 60.07±0.07 59.76±0.22 59.64±0.30 62.71±0.27 64.57±0.22
19,000 66.39±0.26 59.69±0.16 66.82±0.33 66.19±0.19 64.77±0.32 65.59±0.27 64.65±0.23 64.93±0.23 66.89±0.28 67.60±0.08
23,000 68.92±0.15 61.99±0.41 70.54±0.21 69.62±0.32 68.49±0.06 69.03±0.20 68.47±0.50 67.65±0.14 70.81±0.35 70.21±0.32
ImageNet-100 dataset
$|{\mathbb{L}}|$ Random Prob Cover BALD Coreset Uncertainty Entropy DBAL Margin DCoM
400 3.10±0.07 5.28±0.08 1.18±0.21 3.10±0.07 1.19±0.07 1.17±0.13 1.29±0.06 2.22±0.50 5.20±0.02
800 4.61±0.10 7.56±0.15 2.90±0.29 3.68±0.12 2.06±0.24 2.20±0.19 2.15±0.20 3.85±0.34 8.39±0.07
1,200 5.88±0.25 10.10±0.08 4.32±0.33 4.84±0.14 3.42±0.19 3.71±0.26 3.42±0.16 5.54±0.23 11.56±0.07
2,000 9.51±0.33 15.25±0.26 6.73±0.26 6.78±0.15 5.72±0.41 6.15±0.04 5.72±0.15 9.17±0.34 16.27±0.08
6,000 25.67±0.16 32.89±0.12 19.85±0.26 17.98±0.10 20.12±0.11 19.89±0.35 18.72±0.08 26.03±0.08 32.76±0.13
10,000 37.49±0.26 42.00±0.16 31.36±0.46 28.31±0.10 32.13±0.16 31.57±0.28 31.00±0.59 37.33±0.08 41.97±0.44
15,000 46.22±0.30 48.89±0.09 41.16±0.49 38.53±0.18 42.53±0.29 41.88±0.31 41.42±0.13 46.37±0.17 49.47±0.27
20,000 52.20±0.19 51.70±0.32 48.79±0.58 46.59±0.15 49.50±0.25 48.07±0.36 47.89±0.13 53.01±0.19 54.25±0.26
40,000 64.57±0.17 57.60±0.05 64.02±0.37 61.78±0.30 63.94±0.27 63.14±0.37 63.21±0.13 65.19±0.24 65.69±0.17
ImageNet-200 dataset
Table 6: Model accuracy for different sizes of labeled set ${\mathbb{L}}$ and AL strategies
$|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin DCoM
10 15.73±0.90 17.26±0.78 14.92±0.98 12.22±0.91 15.58±1.03 11.88±0.83 11.67±0.48 11.79±1.06 12.81±0.65 16.72±0.94
20 16.68±0.77 22.81±0.28 18.24±0.46 15.28±0.98 18.25±0.97 16.05±0.94 18.48±0.67 18.19±0.63 15.65±1.35 23.47±0.75
30 19.78±0.79 24.91±0.49 19.80±0.38 18.37±1.06 19.41±0.92 17.75±1.05 19.33±0.74 20.35±0.90 18.49±1.16 26.72±0.79
40 21.49±0.48 26.56±0.48 21.57±0.48 21.25±0.95 20.45±1.00 20.86±0.77 21.45±0.74 21.70±0.88 20.97±0.71 28.58±0.69
50 22.19±0.33 28.30±0.42 24.40±0.45 22.13±0.56 21.76±1.07 22.11±0.81 22.75±0.60 22.76±0.91 23.57±0.59 29.20±0.87
200 36.04±0.59 40.91±0.33 35.50±0.57 32.89±0.50 33.22±0.55 33.44±0.76 34.38±0.58 34.95±0.46 36.05±0.53 41.74±0.49
400 44.07±0.60 49.52±0.23 44.18±0.54 43.33±0.45 41.83±0.39 42.39±0.64 42.51±0.46 43.67±0.47 44.35±0.52 47.02±0.73
600 50.02±0.44 52.65±0.35 50.10±0.44 49.56±0.32 48.91±0.32 49.46±0.48 48.85±0.38 49.23±0.49 49.27±0.59 51.54±0.86
800 53.69±0.36 54.45±0.56 54.57±0.63 54.20±0.51 54.42±0.21 53.48±0.38 53.70±0.38 53.98±0.31 54.05±0.56 55.05±0.71
1,400 62.00±0.26 61.54±0.30 63.38±0.25 62.82±0.51 62.72±0.42 62.42±0.27 62.24±0.23 63.04±0.23 63.11±0.25 63.04±0.57
2,600 70.18±0.27 70.50±0.08 72.76±0.19 72.72±0.26 71.41±0.13 72.53±0.23 72.35±0.18 72.97±0.18 72.12±0.40 72.93±0.33
3,800 75.51±0.21 72.97±0.17 76.75±0.15 76.69±0.11 75.75±0.09 76.80±0.21 77.05±0.15 76.85±0.09 76.91±0.10 76.92±0.18
STL-10 dataset
$|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin LDM-s DCoM
10 11.36±0.50 11.80±0.08 13.26±0.79 13.84±1.64 11.36±0.50 11.41±0.85 11.06±1.38 11.66±0.56 10.80±0.66 - 12.62±0.60
20 11.05±0.41 10.96±0.19 12.19±0.67 12.74±1.26 10.99±0.59 11.81±0.84 10.86±0.58 14.94±0.36 12.50±0.95 - 13.30±0.26
30 11.16±0.36 12.19±0.52 12.12±0.51 14.22±2.01 11.10±0.29 12.66±1.28 11.86±0.65 13.50±0.46 11.56±0.79 - 13.91±0.32
40 11.85±0.46 12.11±0.20 12.71±0.70 13.98±2.06 11.57±0.57 14.16±1.63 12.28±0.47 13.10±0.41 11.76±0.75 - 14.32±0.14
50 12.55±0.49 12.86±0.36 12.47±0.89 13.18±2.00 11.76±0.38 14.12±1.00 12.72±0.54 12.88±0.54 12.04±0.68 - 15.09±0.25
200 22.32±0.84 24.34±0.44 20.18±1.09 19.00±1.66 17.42±1.03 21.99±1.11 20.38±1.08 20.98±0.99 22.03±0.87 - 24.58±1.10
400 36.24±1.24 40.42±0.67 35.13±0.40 28.74±2.78 28.88±1.47 29.58±0.55 33.01±1.36 33.06±1.82 34.01±0.18 - 46.90±1.56
600 47.29±0.98 52.67±0.75 46.46±0.87 37.26±1.56 40.24±1.24 42.65±0.68 42.89±1.38 43.99±0.46 46.69±0.50 - 58.34±1.25
800 56.32±0.50 62.96±0.51 56.48±1.33 47.47±2.46 49.74±1.42 50.62±2.03 53.17±1.64 46.45±1.14 55.27±1.05 - 65.90±0.97
1,400 74.00±0.38 73.16±0.21 75.53±0.77 65.33±3.30 68.44±0.69 72.13±0.98 72.81±1.09 73.14±0.69 74.53±0.74 79.02±0.01 78.36±0.36
2,600 83.24±0.48 81.94±0.08 86.56±0.24 83.33±0.96 83.83±0.39 85.80±0.31 84.72±0.54 84.83±0.59 86.33±0.13 83.76±0.00 86.31±0.26
3,800 87.18±0.14 85.34±0.11 89.61±0.25 88.91±0.54 88.86±0.14 89.64±0.09 88.79±0.14 89.36±0.15 89.34±0.18 86.08±0.00 89.28±0.13
5,000 88.79±0.15 87.58±0.20 91.72±0.15 90.93±0.14 91.24±0.17 91.55±0.04 91.13±0.11 91.28±0.12 91.75±0.17 87.58±0.00 91.50±0.08
6,200 90.16±0.04 88.55±0.10 92.64±0.09 92.46±0.04 92.35±0.03 92.63±0.08 92.56±0.16 92.65±0.10 92.55±0.12 88.64±0.00 92.54±0.12
7,400 90.99±0.11 89.71±0.08 93.56±0.05 93.26±0.10 93.02±0.14 93.33±0.05 93.38±0.07 93.38±0.11 93.43±0.13 89.56±0.00 93.47±0.09
SVHN dataset

### v3

@@
-DCoM: Active Learning for All Learners
+DCoM: Competence-Driven Adaptive Active Learning
 Abstract
-Deep Active Learning (AL) techniques can be effective in reducing annotation costs for training deep models.
+Deep Active Learning techniques can be effective in reducing annotation costs for training deep models.
 However, their effectiveness in low- and high-budget scenarios seems to require different strategies, and achieving optimal results across varying budget scenarios remains a challenge.
 In this study, we introduce Dynamic Coverage & Margin mix (DCoM), a novel active learning approach designed to bridge this gap.
-Unlike existing strategies, DCoM dynamically adjusts its strategy, considering the competence of the current model.
+Unlike existing strategies, DCoM dynamically adjusts its strategy, while taking into account the competence of the current model.
 Through theoretical analysis and empirical evaluations on diverse datasets, including challenging computer vision tasks, we demonstrate DCoM’s ability to overcome the cold start problem and consistently improve results across different budgetary constraints.
@@
 1 Introduction
-Deep Learning (DL) algorithms require a lot of data to achieve optimal results.
-In certain scenarios, there is an abundance of unlabeled data but limited capacity for labeling it.
+Deep Learning (DL) algorithms require large amounts of data to achieve optimal results.
+In some situations, there is an abundance of unlabeled data but limited capacity to label it.
 In fields such as medical imaging, an invaluable resource — doctors themselves — serves as a costly oracle.
-Active learning algorithms aim to address this challenge by reducing the burden of labeling and making it more effective.
-Unlike traditional supervised learning frameworks, Active Learning (AL) can affect the construction of the labeled set, possibly by leveraging knowledge about the current learner.
+Active learning (AL) algorithms aim to address this challenge by reducing the labeling burden and enhancing its effectiveness (see illustration in Fig.
+1).
+Unlike traditional supervised learning frameworks, AL can influence the construction of the labeled set, potentially by leveraging knowledge about the current learner.
 Accordingly, the initial goal of AL is to select $q$ examples to be annotated, where $q$ represents the number of examples that can be sent to the oracle.
-AL has already demonstrated tangible contributions across various domains such as computer vision tasks [40], NLP [31, 20] and medical imaging [10].
+AL has already demonstrated tangible contributions across various domains such as computer vision tasks (e.g.
+Yuan et al.
+2023), NLP (e.g.
+Siddhant and Lipton 2018; Kaseb and Farouk 2023) and medical imaging (e.g.
+El-Hasnony et al.
+2022).
 These examples highlight the importance of advancing AL to achieve even greater impact.
-The choice of an AL strategy depends on both the learner’s inductive biases and the nature of the problem.
-Recent research suggests that the optimal active learning strategy varies with the budget size, where budget refers to the size of the training set.
-When the budget is large, methods based on uncertainty and diversity sampling are most effective.
-Conversely, when the budget is small, methods centered on typicality and diversity are more appropriate.
-However, there is no single active learning strategy that is suitable for all budget regimes.
-Our study aims to address this challenge by dynamic selection of the best examples based on the current state of the learner and the current budget.
-As briefly reviewed below, recent research often categorizes active sampling methods according to the budget regime.
-In contrast, our study takes a different approach by recognizing the need to adapt sample selection methods based on the learner’s competence.
+Figure 1: Illustration of the AL iterative process: DCoM introduces an additional step to estimate the current learner’s competence, which is then utilized during the active sampling phase.
+The choice of an active learning (AL) strategy depends on both the learner’s inductive biases and the nature of the problem.
+Recent research indicates that the optimal AL strategy varies with the budget size, where the budget refers to the size of the training set.
+For larger budgets, uncertainty and diversity sampling methods are most effective, while for smaller budgets, strategies focused on typicality and diversity are more suitable.
+However, no single AL strategy is universally optimal across all budget levels.
+Our study addresses this challenge by dynamically selecting the best examples based on the learner’s current state and the available budget.
+Recent research, as briefly reviewed below, often categorizes active sampling methods based on budget regimes.
+In contrast, our study takes a different approach by emphasizing the need to adapt sample selection methods according to the learner’s competence.
 We aim to develop a single algorithm that dynamically adjusts its sampling strategy in response to the learner’s evolving abilities.
-This transition from budget to learner’s competence represents a significant advancement in active learning research, as it allows us to move beyond rigid budget allocations and towards a more adaptive approach that optimizes performance based on the learner’s abilities.
-In this work, we introduce a novel method called Dynamic Coverage & Margin mix (DCoM), which dynamically adjusts its selection strategy to deliver optimal results.
-We begin by examining a simplified theoretical framework (Section 2), where we integrate a typicality test with an uncertainty test based on the budget and its coverage.
-Inspired by the insights gained from this analysis, we introduce DCoM (Section 3), which combines these characteristics.
-DCoM aims to offer a flexible solution for any budget by selecting the most appropriate examples that are most suitable to the current budget.
-We validate DCoM through an extensive empirical study using various vision datasets (Section 4).
-Our findings indicate that DCoM effectively achieves superior performance across all budget ranges.
+This shift from a budget-focused to a competence-driven approach represents a significant advancement in active learning research, allowing for a more adaptive strategy that optimizes performance based on measurable learner’s capabilities.
+In this work, we introduce a novel method called Dynamic Coverage & Margin Mix (DCoM), which dynamically adjusts its selection strategy by minimizing a new loss function that incorporates a measure of competence.
+We start by exploring a simplified theoretical framework (Section 2), where we integrate a typicality test with an uncertainty test based on the budget and a notion of coverage.
+Building on the insights from this analysis, we introduce DCoM (Section 3), which combines these elements to provide a flexible solution for any budget.
+Thus, DCoM selects the most appropriate examples suited to the current budget.
+We validate DCoM through an extensive empirical study using various vision datasets (Section 4), demonstrating that DCoM consistently achieves superior performance across all budget ranges.
 Relation to prior art.
-Over the past years, active learning has been an active area of research [34, 29, 7, 32].
-The realm of Active Learning (AL) is segmented into three problem settings: membership query synthesis, stream-based selective sampling, and pool-based active learning [29].
-The majority of current AL research adheres to the pool-based AL setting, where samples are drawn from a vast pool of unlabeled data and annotations are solicited from an oracle.
+Over the past years, active learning has been an active area of research (Wang et al.
+2023; Settles 2009; Cohn, Ghahramani, and Jordan 1996; Tharwat and Schenck 2023; Xie et al.
+2023; Schmidt et al.
+2024).
+The realm of Active Learning (AL) is segmented into three problem settings: membership query synthesis, stream-based selective sampling, and pool-based active learning (Settles 2009).
+Much of the recent AL work follows the pool-based setting, where samples are selected from a large pool of unlabeled data and annotations are obtained from an oracle.
 In this setting, AL algorithms operate within a given budget ($\mathit{b}$), selecting a subset of unlabeled examples to send to an oracle for labeling.
@@
 In discussions of pool-based AL, two key characteristics that play a crucial role during the active sampling process are uncertainty and diversity.
-Some AL algorithms focus solely on uncertainty sampling [23, 30, 4, 43, 18, 12, 27, 37, 11], while others prioritize diversity sampling [28, 19, 39, 15].
-Other algorithms combine both uncertainty and diversity in their approaches [38, 2, 36, 21, 5, 1].
-Over time, the cold start problem emerged, highlighting the challenges faced by AL algorithms when operating with a small budget.
-It seems that uncertainty and diversity do not capture this difficulty.
-Chen et al.
-[2], for example, describe the cold start problem as a combination of two issues: unbalanced sampling across different classes (which relates to diversity) and unbalanced sampling within each class, where uncertain examples are favored over certain ones at the outset.
-In low-budget scenarios, typical examples often offer the best outcomes for the model [2, 39, 15, 14].
+Some AL algorithms focus solely on uncertainty sampling (Lewis and Gale 1994; Shannon 1948; Cho et al.
+2023; Zhao et al.
+2021; Houlsby et al.
+2011; Gissin and Shalev-Shwartz 2019; Ranganathan et al.
+2017; Woo 2022; Gal, Islam, and Ghahramani 2017; Parvaneh et al.
+2022; Jung, Kim, and Lee 2023), while others prioritize diversity sampling (Sener and Savarese 2018; Hu, Mac Namee, and Delany 2010; Yehuda et al.
+2022; Hacohen, Dekel, and Weinshall 2022).
+Other algorithms combine both uncertainty and diversity in their approaches (Yang et al.
+2015; Chen et al.
+2024; Wen, Pizarro, and Williams 2023; Kirsch, Van Amersfoort, and Gal 2019; Chowdhury, Hamerly, and McGarrity 2024; Ash et al.
+2019; Kim et al.
+2021).
+Over time, the cold start problem has emerged, highlighting the difficulties faced by AL algorithms when operating with a small budget.
+It seems that uncertainty and diversity are ill-suited to address this challenge.
+Accordingly, Chen et al.
+(2024) describe the cold start problem as a combination of two issues: unbalanced sampling across different classes (which relates to diversity).
+and unbalanced sampling within each class, where uncertain examples are favored over certain ones at the outset.
+In low-budget scenarios, typical examples often offer the highest benefits for training the model (Hacohen, Dekel, and Weinshall 2022; Yehuda et al.
+2022; Chen et al.
+2024).
 This implies that in such contexts, AL algorithms should aim to maximize both certainty and diversity.
 Conversely, in high-budget scenarios, the emphasis shifts towards maximizing uncertainty while maintaining diversity.
-[14] proposes a method to select the most suitable approach based on the current state.
-This method involves comparing multiple active learning strategies during runtime, rather than recommending a single universal approach.
+While (Hacohen and Weinshall 2024) proposes a method to select the most suitable approach for the current state, the method involves the comparison of multiple active learning strategies at each state, rather than recommending a single universal approach.
 Summary of contribution.
 (i) Expand a theoretical framework to analyze Active Learning (AL) strategies in embedding spaces.
-(ii) Introduce DCoM, a multi-budget strategy that significantly outperforms previous methods in low and medium budget scenarios and matches their performance in high budget scenarios.
-(iii) Initiates a transition from a budget-based approach to the learner’s competence, introducing coverage as a measure designed to predict this competence.
+(ii) Introduce DCoM, a universal strategy that significantly outperforms previous methods in low and medium budget scenarios, and matches them in high budget scenarios.
+(iii) Initiate a transition from a budget-based discussion to the learner’s competence, while introducing coverage as a measure designed to predict competence.
 2 Theoretical Framework
-The framework of active learning adopted here can be formalized as follows: Let ${\mathbb{X}}$ denote the input domain, ${\mathbb{Y}}$ the target domain, and $\mathcal{L}:{\mathbb{X}}\to{\mathbb{Y}}$ the target learner.
+Let $\mathbb{X}$} denote the input domain, $\mathbb{Y}$} the target domain, and $\mathcal{L}:{\mathbb{X}}\to{\mathbb{Y}}$ the trained learner.
 Before seeing any labels, $\mathcal{L}$ may be a random feasible hypothesis, or it may be initialized by either transfer learning or self training.
+Let $b$ denote the size of the labeled training set, aka budget.
 Active learning may involve a single step, or an iterative process with repeated active learning steps.
-In each step of active a learning, the learner actively seeks labels by choosing a set of unlabeled points as a query set, to be labeled by an oracle/teacher.
-Subsequently, this set of labeled points is added to the learner’s supervised training set, and the learner is retrained or fine-tuned.
-As discussed in the introduction, it has been shown that different active learning strategies are suitable for different budgets $b$.
-Henceforth, $b$ refers to the number of labeled examples known to the learner at the beginning of an active learning step.
+In each AL step, the learner actively seeks labels by choosing a set of unlabeled points as a query set, to be labeled by an oracle/teacher.
+Subsequently, this set of labeled points is added to the learner’s supervised training set, and the learner is either retrained or fine-tuned.
+As discussed in Section 1, it has been shown that different active learning strategies are suited for different budgets $b$.
 With a low budget $b$, strategies that do not rely on the outcome of the learner are most suitable.
-When $b$ is high, it is beneficial to take into consideration the confidence of the learner when choosing an effective query set.
-What makes a budget high or low was left vague in previous analysis, as it clearly depends on the specific application and dataset.
+When $b$ is high, it is beneficial to consider the confidence of the learner when choosing an effective query set.
+What makes a certain budget $b$ high or low is left vague in previous work, as it depends on both the hypothesis space (or architecture) and the dataset.
 To achieve an effective active learning protocol suitable for all learners, irrespective of budget, we aim to devise a universal objective function, whose minimization is used to select the query set.
-Next, we discuss its envisioned form for selecting the query set at the beginning of each AL step.
-We begin our discussion by proposing to replace the notion of budget with the notion of competence: a network is competent if its generalization score is relatively high, and vice versa.
-This proposal is more formally justified in Section 2.3.
-We then rephrase the intuition stated above as follows: When the learner’s competence is low, the objective function should prioritize typicality and diversity of selected queries (see discussion in the introduction) irrespective of the learner’s predictions.
+We begin by proposing to replace the notion of budget with the notion of competence, designed to track the generalization ability of $\mathcal{L}$ when trained with budget $b$.
+The notion of competence, we hypothesize, is more universal and less dependent on the type of learner and specific dataset (see discussion in Section 2.3).
+We then rephrase the intuition stated above (see Section 1) as follows: When the learner’s competence is low, the objective function should prioritize typicality and diversity of selected queries irrespective of the learner’s predictions.
 When the learner’s competence is high, its uncertainty in prediction should be given high priority.
@@
 ${\mathcal{O}}(x)=(1-{S_{\mathcal{L}}})\cdot{{{\mathcal{O}_{low}}}(x)}+{S_{\mathcal{L}}}\cdot{{{\mathcal{O}_{high}}}(x)}$ (1)
-In Section 2.2 we discuss the design of ${\mathcal{O}_{low}}(x)$.
-In order to tackle the challenge of active learning when the competence of $\mathcal{L}$ is low, we utilize the point coverage framework introduced in [39].
+After the introduction of necessary preliminaries in Section 2.1 we discuss in Section 2.2 the design of ${\mathcal{O}_{low}}(x)$.
+We begin with the point coverage framework introduced by Yehuda et al.
+(2022).
 This approach relies on self-supervised data representation, which is blind to the learner’s performance.
-We refine their analysis and obtain an improved generalization bound for the 1-Nearest Neighbor (1-NN) classification model that depends on the local geometry of the data.
-Beforehand, in Section 2.1 we discuss the necessary preliminaries required to introduce the coverage approach.
-In Section 2.3 we discuss our choice of ${S_{\mathcal{L}}}$, and connect it to the notion of budget discussed in previous work.
-We select ${\mathcal{O}_{high}}(x)$ to equal one minus the normalized lowest response between the two highest softmax outputs of $\mathcal{L}$ at $x$, or Margin, as this is a common measure of prediction uncertainty.
+We refine their analysis and obtain an improved generalization bound for the Nearest Neighbor classification model that depends on the local geometry of the data.
+Next, in Section 2.3 we discuss our choice of $S_{\mathcal{L}}$}, and connect it to the notion of budget discussed in previous work.
+Finally, we select ${\mathcal{O}_{high}}(x)$ to equal one minus the normalized lowest response between the two highest softmax outputs of $\mathcal{L}$ at $x$, or Margin, as this is a common uncertainty measure.
 2.1 Preliminaries
@@
 $\pi\left(\delta\right)=P\left(\left\{x\in{\mathbb{X}}:B_{\delta}\left(x\right)\text{ is pure}\right\}\right).$
-Notice that $\pi(\delta)$ is monotonically decreasing, as can be readily verified.
-In [39] it is shown that the generalization error of the 1-NN classifier $f^{N}$ is bounded as follows
-${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\left(1-P(C({\mathbb{L}},\delta))\right)+(1-\pi(\delta)).$ (2)
-Subsequently, an algorithm is proposed that minimizes the first term in (2) by maximizing the coverage probability $P(C({\mathbb{L}},\delta)$, while ignoring the second term that is assumed to be fixed.
+Note that $\pi(\delta)$ is monotonically decreasing, as can be readily verified.
+In (Yehuda et al. 2022) it is shown that the generalization error of the 1-NN classifier $f^{N}$ is bounded as follows
+${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\left(1-P[C({\mathbb{L}},\delta)]\right)+(1-\pi(\delta)).$ (2)
+Subsequently, an algorithm is proposed that minimizes the first term in (2) by maximizing the coverage probability $P[C({\mathbb{L}},\delta)]$, while ignoring the second term that is assumed to be fixed.
 Below, we begin by refining the bound, focusing more closely on its second term $(1-\pi(\delta))$.
@@
 2.2 Refined error bound
-Define the indicator random variable $I_{\delta}(x)=\mathbbm{1}_{\{x\in{\mathbb{X}}:B_{\delta}(x)\text{ is pure}\}}$. By definition,
+Define the indicator random variable $I_{\delta}(x)$ as follows:
+$I_{\delta}(x)={\mathds{1}_{\{x\in{\mathbb{X}}:B_{\delta}(x)\text{ is pure}\}}}.$
+By definition,
 $\pi\left(\delta\right)=P\left(\left\{x\in{\mathbb{X}}:B_{\delta}\left(x\right)\text{ is pure}\right\}\right)={\mathbb{E}}[I_{\delta}(x)].$
 Since the distribution of $\mathbb{X}$} is not known apriori, we use the empirical distribution to approximate the expected value of this random variable.
-In accordance, given a labeled set ${\mathbb{L}}=\left\{x_{i}\right\}_{i=1}^{b}$:
-$\begin{split}\hat{\pi}(\delta)&=\hat{{\mathbb{E}}}[\mathbbm{1}_{\{x\in{\mathbb{L}}:B_{\delta}(x)\text{ is pure}\}}]=\frac{1}{b}{\mathbb{E}}[\sum_{i=1}^{b}\mathbbm{1}_{\{B_{\delta}(x_{i})\text{ is pure}\}}]=\frac{1}{b}\sum_{i=1}^{b}{\mathbb{E}}[\mathbbm{1}_{\{B_{\delta}(x_{i})\text{ is pure}\}}]\\ &=\frac{1}{b}\sum_{i=1}^{b}P(B_{\delta}(x_{i})\text{ is pure}).\end{split}$ (3)
+Thus, with labeled set ${\mathbb{L}}=\left\{x_{i}\right\}_{i=1}^{b}$
+[…]
 With this approximation
-${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big{[}1-P(C({\mathbb{L}},\delta))\Big{]}+\Big{[}1-\frac{1}{b}\sum_{i=1}^{b}P(B_{\delta}(x_{i})\text{ is pure})\Big{]}+\varepsilon.$ (4)
-where $\varepsilon$ bounds the error of the empirical approximation in (3).
+$\begin{split}{\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big[1-P[C({\mathbb{L}},\delta)]\Big]\qquad\qquad\\ +\Big[1-\frac{1}{b}\sum_{i=1}^{b}P[B_{\delta}(x_{i})\text{ is pure}]\Big]+\varepsilon.\end{split}$ (3)
+where $\varepsilon\underset{b\to\infty}{\to}0$ bounds the error of the approximation $\hat{\pi}(\delta)$.
 Note that the refined bound in (3) depends on the purity separately at each labeled point in $\mathbb{L}$}.
-We further refine this bound by allowing the fixed raidus $\delta$ to be chosen individually, where $\delta_{i}$ denotes the radius of point $x_{i}\in{\mathbb{L}}$.
+Next, we further refine this bound by allowing the fixed radius $\delta$ to be chosen individually, where $\delta_{i}$ denotes the radius of point $x_{i}\in{\mathbb{L}}$.
 Let $\Delta=[\delta_{i}]_{i=1}^{b}$ denote the list of individual radii corresponding to ${\mathbb{L}}=\{x_{i}\}_{i=1}^{b}$.
 The cover defined by $\Delta$ is $C({\mathbb{L}},\Delta)=\bigcup_{(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta}B_{\delta_{i}}(x_{i})$.
-We consider the normalized Nearest Neighbor (nNN) classification algorithm, where the nearest neighbor is chosen by solving $f^{N}(x))=\text{argmin}_{x_{i}\in{\mathbb{L}}}\frac{d(x,x_{i})}{\delta_{i}}$.
+With this choice, it is necessary to replace the 1-NN classifier with a suitable variant of nearest neighbor classification - the normalized Nearest Neighbor (1-nNN) classifier.
+In 1-nNN, distances to point $x_{i}\in{\mathbb{L}}$ are normalized by the corresponding $\delta_{i}$, after which the nearest neighbor $f^{N}(x)=\text{argmin}_{x_{i}\in{\mathbb{L}}}\frac{d(x,x_{i})}{\delta_{i}}$ is computed.
+This transition preserves the original problem settings, and guarantees that if a labeled point $c\in{\mathbb{L}}$ assigns labels to other points, it also covers them (see Appendix B).
+Consequently, if a point $x$ is mislabeled, there exists a $c\in{\mathbb{L}}$ that covers it, leading to a reduction in purity.
 The bound on the error in (3) becomes
-${\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big{[}1-P(C({\mathbb{L}},\boldsymbol{\Delta}))\Big{]}+\Big{[}1-\frac{1}{b}\sum_{i=1}^{b}P(B_{\delta_{i}}(x_{i})\text{ is pure})\Big{]}+\varepsilon.$ (5)
-2.3 Competence score $\boldsymbol{{S_{\mathcal{L}}}}$
-In order to obtain a useful competence score [^1] When the labeled set ${\mathbb{L}}$ is large, it may be possible to set aside a validation set to directly estimate the competence of learner $\mathcal{L}$, but this is not feasible when the labeled set ${\mathbb{L}}$ is small., which can be effectively used in objective function (1), we require that it satisfies the following conditions:
-  1. (i)
-    Lie in the range $[0,1]$.
-  2. (ii)
-    Use all the data.
-  3. (iii)
-    Monotonically increase with the competence of learner $\mathcal{L}$.
-In accordance, we propose to use the probability of coverage:
-${S_{\mathcal{L}}}({\mathbb{L}},\Delta)=\frac{1+e^{-k(1-a)}}{1+e^{-k(P(C({\mathbb{L}},\Delta))-a)}}$ (6)
-This function follows a logistic growth curve, with parameters defined as follows: $a\in(0,1)$ is the midpoint and $k\in(0,\infty)$ determines the curve’s steepness.
-First, we note that this score satisfies the conditions stated above for the nNN classifier: (i)-(ii), as it is a function in $[0,1]$ that depends on all the data ${\mathbb{X}}$, while (iii) follows from the error bound (5).
-Below, we prove that it is also monotonically increasing with budget $b$.
-More specifically, prior work showed empirically and theoretically that the training budget $b=|{\mathbb{L}}|$ provides a useful clue for the desirable trade-off between ${{\mathcal{O}_{low}}}(x)$ and ${{\mathcal{O}_{high}}}(x)$ in (1), which is captured by score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ defined in (6).
-We next show that ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is monotonically increasing with $b$.
-This implies that definition (6) agrees with empirical evidence in prior work.
+$\begin{split}{\mathbb{E}}\left[f^{N}(x)\neq f(x)\right]\leq\Big[1-P[C({\mathbb{L}},\Delta)]\Big]\qquad\qquad\\ +\Big[1-\frac{1}{b}\sum_{i=1}^{b}P[B_{\delta_{i}}(x_{i})\text{ is pure}]\Big]+\varepsilon.\end{split}$ (4)
+This bound describes a trade-off between probability coverage and purity, and guides the algorithm design in Section 3.
+Note that in (4), the cover $P[B_{\delta_{i}}(x_{i})\text{ is pure}]$ is refined locally, using $B_{\delta_{i}}$ instead of $B_{\delta}$.
+[…]
+In order to obtain a useful competence score [^1] When the labeled set $\mathbb{L}$} is large, it may be possible to set aside a validation set to directly estimate the competence of learner $\mathcal{L}$, but this is not feasible when the labeled set $\mathbb{L}$} is small., which can be effectively used in objective function (1), we require that it satisfies the following conditions: (i) Lie in the range $[0,1]$; (ii) Depend on the entire dataset; (iii) Monotonically increase with the competence of learner $\mathcal{L}$.
+In accordance, we propose to use a constrained Sigmoid function that depends on the probability of coverage:
+${S_{\mathcal{L}}}({\mathbb{L}},\Delta)=\frac{1+e^{-k(1-a)}}{1+e^{-k(P[C({\mathbb{L}},\Delta)]-a)}}$ (5)
+This score follows the sigmoid curve, where $a$ denotes its midpoint and $k$ the steepness of the curve.
+In Appendix C, we analyze the choice of these parameters and its impact on the algorithm.
+Note that this score satisfies the conditions stated above for the nNN classifier: (i)-(ii), as it is a function in $[0,1]$ that depends on $P[C({\mathbb{L}},\Delta)]=\frac{|C({\mathbb{L}},\Delta)|}{|{\mathbb{X}}|}$, while (iii) follows from the error bound (4).
+Below, we demonstrate that $S_{\mathcal{L}}$} is monotonically increasing with the budget $b$, which supports the alignment of definition (5) with empirical evidence from prior work.
+Specifically, Hacohen, Dekel, and Weinshall (2022) showed both empirically and theoretically that the training budget $b=|{\mathbb{L}}|$ serves as a valuable indicator for the relative suitability of ${{\mathcal{O}_{low}}}(x)$ and ${{\mathcal{O}_{high}}}(x)$, which are employed in (1).
+In our formulation, this relationship informs a trade-off represented by the score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ defined in (5).
+It remains to be shown that ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is indeed monotonically increasing with $b$.
 Proposition.
@@
 First, we note that
-${\mathbb{L}}\subseteq{\mathbb{L}}^{\prime}\implies C({\mathbb{L}},\Delta)=\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\subseteq\left(\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)\cup\left(\bigcup_{x_{i}\in{\mathbb{L}}^{\prime}\setminus{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)=C({\mathbb{L}}^{\prime},\Delta^{\prime})$
-Since probability is monotonic, it follows that $P(C({\mathbb{L}},\Delta))\leq P(C({\mathbb{L}}^{\prime},\Delta^{\prime}))$.
+$\begin{split}{\mathbb{L}}\subseteq{\mathbb{L}}^{\prime}&\implies C({\mathbb{L}},\Delta)=\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\\ &\subseteq\left(\bigcup_{x_{i}\in{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)\cup\left(\bigcup_{x_{i}\in{\mathbb{L}}^{\prime}\setminus{\mathbb{L}}}B_{\delta_{i}}(x_{i})\right)\\ &=C({\mathbb{L}}^{\prime},\Delta^{\prime})\end{split}$
+Since probability is monotonic, it follows that $P[C({\mathbb{L}},\Delta)]\leq P[C({\mathbb{L}}^{\prime},\Delta^{\prime})]$.
 From (5), and by the definition of the logistic function using the provided parameters limits, ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is monotonically increasing in $\mathbb{L}$}, which implies that ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)\leq{S_{\mathcal{L}}}({\mathbb{L}}^{\prime},\Delta^{\prime})$.
@@
 3 Our method: DCoM
-The error bound in (5) guides our method for active query selection, which aims to minimize this bound by maximizing the coverage $P(C({\mathbb{L}},\Delta))$ and individual purity $\pi(\delta_{i})=P(B_{\delta_{i}}(x_{i})\text{ is pure})$.
-Recall that ${\mathbb{L}}$ denotes the set of labeled points, $\delta_{i}$ the influence radius of $x_{i}\in{\mathbb{L}}$, $\Delta=[\delta_{i}]_{i=1}^{b}$, and
-$C({\mathbb{L}},\Delta)=\bigcup_{(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta}B_{\delta_{i}}(x_{i}),\qquad\qquad\pi(\delta_{i})=P\left(\left\{x\in B_{\delta_{i}}\left(x_{i}\right):f(x)=f(x_{i})\right\}\right)$
-Since it was shown in [39] that maximizing $P(C({\mathbb{L}},\delta))$ is already NP-hard, we opt for developing a greedy algorithm for actively selecting the query set, to be termed DCoM.
+[…]
+(a) Initial active sampling using $\delta_{0}$ in DCoM, following ProbCover approach.
+(b) DCoM’s updated $\Delta$ list, which includes a suitable $\delta$ value for each example in $\mathbb{L}$}
+Figure 2: Illustration of Step 4: $\Delta$-Adjustment in DCoM, comparing the coverage obtained by DCoM (b) with ProbCover (a) using budget $b=15$.
+Selected points are marked by x, and the covered area is shaded in light gray.
+While ProbCover maintains a constant $\delta$ for each example, DCoM seeks to maximize coverage while preserving the purity of the balls by fitting a suitable radius for each center.
+3.1 Framework and definitions
+$\forall x\in{\mathbb{X}}$, let $M(x)$ denote the normalized margin between the two highest softmax outputs from the last trained model $\mathbb{M}$}.
+Define $\Delta$ as the set of individual $\delta$-values for each labeled example in $\mathbb{L}$}.
+As in the theoretical analysis, $P[C({\mathbb{L}},\Delta)]$ denotes the probability of coverage for $\mathbb{L}$} given $\Delta$.
+The competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is defined by the objective function (5).
+The graph $\{G_{\delta}=(V,E)\}$ is a directed adjacency graph where nodes are samples, and edges connect pairs of nodes if their distance in the embedding space is less than $\delta$.
 3.2 Active learning method
-The algorithm involves several steps:
-Initialization: Define a suitable embedding space for the data and initialize essential parameters.
-If ${\mathbb{L}}\neq\emptyset$, train model ${\mathbb{M}}$ on ${\mathbb{L}}$, and compute $\boldsymbol{\delta}$-expansion on the elements of ${\mathbb{L}}$.
-Otherwise, $\Delta=[\delta_{i}=\delta_{0}]_{i=1}^{b}$.
-Repeat /* iterative active learning steps */
-  1. (i)
-    Active query selection: determine which examples to select for annotation.
-  2. (ii)
-    Model training: obtain labels and train a deep learning model ${\mathbb{M}}$.
-  3. (iii)
-    $\boldsymbol{\delta}$-expansion: use ${\mathbb{M}}$ to update $\delta_{i}$ for the recently selected examples.
-Until the accumulated budget is exhausted.
-We next provide a detailed description of each step.
-Initialization: We begin by seeking a suitable embedding space using ${\mathbb{U}}\cup{\mathbb{L}}$, where distances are intended to be inversely correlated with semantic similarity.
-Self-supervised or representation learning may be used to this end.
-We choose an initial radius $\delta_{0}$ as described in [39].
-(i) Active query selection: In this phase (see Alg 1 below for pseudo-code), the algorithm selects $q$ examples from ${\mathbb{U}}$ to send for labeling, where $q$ denotes the query set size.
-To this end, it first computes several parameters:
-Algorithm 1 DCoM, Active sampling
-1:  Input: unlabeled pool ${\mathbb{U}}$, labeled pool ${\mathbb{L}}$, query size $q$, list of ball sizes $\Delta$, trained model ${\mathbb{M}}$ if ${\mathbb{L}}\neq\emptyset$
-2:  Output: a set of points to query, and the coverage of ${\mathbb{L}}$
-3:  $\forall x\in{\mathbb{U}}$, $M(x)\leftarrow$ normalized margin between the two highest softmax outputs of ${\mathbb{M}}$ if ${\mathbb{L}}\neq\emptyset$, 1 otherwise
-4:  Compute $P(C({\mathbb{L}},\Delta))$
-5:  Compute ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$
-6:  $\delta_{\text{avg}}\leftarrow$ Average($\Delta$)
-7:  $G=(V={\mathbb{U}}\cup{\mathbb{L}},E=\{(x,x^{\prime}):x^{\prime}\in B_{\delta_{\text{avg}}}(x)\})$
-8:  for $(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta$ do
-9:     $\forall x\in B_{\delta_{i}}(x_{i})$: $\forall e=(v,x)\in E$, remove $e$eqparboxCOMMENT# DON’T COVER POINTS TWICE
-10:     $\forall e=(x_{i},v)\in E$, remove $e$
-11:  end for
-12:  ${{\textit{Q}}}\leftarrow\emptyset$
-13:  for $i=1,...,q$ do
-14:     $\forall x\in{\mathbb{U}}$, $\text{ODR}(x)\leftarrow$Out-Degree Rank
-15:     $\forall x\in{\mathbb{U}}$, $R(x)\leftarrow{S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))$ + (1 - ${S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x)$
-16:     $x_{max}\leftarrow\{\text{argmax}_{x\in{\mathbb{U}}}R(x)\}$
-17:     ${\textit{Q}}\leftarrow{{\textit{Q}}}\cup\{x_{max}\}$
-18:     $\forall x\in B_{\delta_{avg}}(x_{max})$: $\forall e=(v,x)\in E$, remove $e$
-19:     $M(x_{max})\leftarrow\text{1}$
-20:  end for
-21:  return  Q, $P(C({\mathbb{L}},\Delta))$
-  • $M(x),~{}\forall x\in{\mathbb{U}}$: the normalized margin between the two highest softmax outputs using the last trained model if ${\mathbb{L}}\neq\emptyset$, 1 otherwise.
-  • $\Delta$: $\Delta=[\delta_{i}=\delta_{0}]_{i=1}^{b}$ if ${\mathbb{L}}=\emptyset$, otherwise use $\Delta$ from the last AL iteration.
-  • $\delta_{\text{avg}}$: the average of $\Delta$.
-  • $P(C({\mathbb{L}},\Delta))$: the probability of coverage of the current ${\mathbb{L}}$ with its corresponding $\Delta$.
-  • ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$: the competence score of objective function (6) if ${\mathbb{L}}\neq\emptyset$, 0 otherwise.
-  • $\{G_{\delta_{\text{avg}}}=(V,E)\}$: a directed adjacency graph where nodes represent samples and edges connect pairs of nodes if the distance between them in the embedding space is smaller than $\delta_{\text{avg}}$.
-In order not to cover points more than once, $\forall x_{i}\in{\mathbb{L}}$, $\forall x\in B_{\delta_{i}}(x_{i})$ and $\forall{v}\in{V}$ – prune incoming edges $e=(v,x)\in E$.
-Additionally, we prune all outgoing edges $e=(x_{i},v)\in E$.
-Next, $q$ points are to be selected from ${\mathbb{U}}$ in a greedy manner as follows:
+Active Learning (AL) is performed iteratively (see Fig.
+1): at each iteration $i$, a query set Q of $q$ unlabeled examples is selected based on a specific strategy, where $q$ denotes the query set size and Q denotes the current query set.
+These examples are then labeled by an oracle, added to the labeled set $\mathbb{L}$}, and removed from the unlabeled pool $\mathbb{U}$}.
+Subsequently, the algorithm parameters may be updated as needed, depending on the specific requirements of the algorithm.
+This process is repeated until the label budget is exhausted or predefined termination conditions are met.
+Our AL method DCoM involves several distinct steps:
+Step 1: Initialization.
+Identify an appropriate embedding space for ${\mathbb{U}}\cup{\mathbb{L}}$, where distances are expected to be inversely correlated to semantic similarity.
+This can be achieved using self-supervised or representation learning techniques.
+Select an initial radius $\delta_{0}$ as outlined in (Yehuda et al.
+2022), and initialize the parameter $\Delta=\emptyset$.
+If $\mathbb{L}$$\neq\emptyset$, skip Step 2 for now and begin the iterations with Step 3.
+Step 2: Active sampling.
+Select $q$ examples from $\mathbb{U}$} for labeling (see Alg 1 for the pseudo-code).
+Begin by computing the normalized margin $M(x)$ $\forall x\in{\mathbb{U}}$ using the latest model $\mathbb{M}$}.
+Next, construct a directed adjacency graph $\{G_{\delta_{\text{avg}}}=(V,E)\}$, where $\delta_{\text{avg}}$ is the average of $\Delta$, or $\delta_{0}$ if $\Delta$ is empty.
+To avoid covering points multiple times, $\forall x_{i}\in{\mathbb{L}}$, $\forall x\in B_{\delta_{i}}(x_{i})$ and $\forall v\in{V}$, prune all incoming edges $e=(v,x)\in E$.
+Additionally, prune all outgoing edges $e=(x_{i},v)\in E$.
+Using this graph, compute the Out-Degree-Rank (ODR) $\forall x\in{\mathbb{U}}$.
+Likewise, compute the probability of coverage $P[C({\mathbb{L}},\Delta)]$ of the current $\mathbb{L}$} and its corresponding $\Delta$, and use these quantities to determine ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$.
+Now select $q$ points from $\mathbb{U}$} in a greedy manner as follows:
   1. 1.
@@
   2. 2.
-    $\forall x\in{\mathbb{U}}$, compute a ranking score $R(x)={S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))+(1-{S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x)$.
+    $\forall x\in{\mathbb{U}}$, compute a ranking score
+    $\hskip-17.07182ptR(x)={S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))+(1-{S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x).$
   3. 3.
-    Select the vertex $x_{\text{max}}$ with the highest score, $x_{\text{max}}=\text{argmax}_{x\in{\mathbb{U}}}R(x)$.
+    Select the vertex $x_{\text{max}}$ with the highest score
+    $x_{\text{max}}=\text{argmax}_{x\in{\mathbb{U}}}R(x).$
   4. 4.
@@
     Set $M(x_{\text{max}})=1$.
-(ii) Model training:
-  1. 1.
-    Obtain labels for query set Q, returned in step (i) of Active sampling.
-  2. 2.
-    Remove set Q from ${\mathbb{U}}$ and add it to ${\mathbb{L}}$.
-  3. 3.
-    Train the given model ${\mathbb{M}}$ using ${\mathbb{L}}$ (and ${\mathbb{U}}$ if the learner is semi-supervised).
-(iii) $\boldsymbol{\delta}$-expansion for new samples.
-In this step (see Alg 2 below for pseudo-code), we compute a suitable $\delta_{i}$ for each new labeled example $x_{i}\in{\textit{Q}}$, and add it to $\Delta$.
-First, we use the updated model $M$ to predict labels for all the points in the unlabeled set ${\mathbb{U}}$.
-We also compute a purity threshold $\tau=m\cdot P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))+b$, where $m$ and $b$ are hyperparameters and $P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))$ the probability of coverage using the old labeled set - has been computed in the first Active sampling step.
-For each new labeled example $v\in{\textit{Q}}$, we search for the largest radius $\delta_{\text{opt}}$ such that its purity $P\left(\left\{x\in B_{\delta_{\text{opt}}}\left(v\right):f(x)=f(v)\right\}\right)$ is still larger than $\tau$.
-Assuming that $\pi(B_{\delta}(x))$ a monotonic with $\delta$, we can leverage binary search.
-Algorithm 2 DCoM, $\boldsymbol{\delta}$-expansion
-1:  Input: unlabeled pool ${\mathbb{U}}$, labeled pool ${\mathbb{L}}$, query pool Q, maximal ball size $\delta_{\text{max}}$, list of ball sizes $\Delta$ for ${\mathbb{L}}\setminus{{\textit{Q}}}$, trained model ${\mathbb{M}}$ from the current iteration and $P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))$ from Active sampling
-2:  Output: updated list of ball sizes $\Delta$
-3:  $\tau\leftarrow m\cdot P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))+b$
-4:  $\hat{{\mathbb{Y}}}\leftarrow$ The predicted label for each $x\in{\mathbb{U}}$ using the model ${\mathbb{M}}$
-5:  for $v\in{{\textit{Q}}}$ do
-6:     $\delta_{\text{opt}}\leftarrow\text{argmax}_{\delta}[P\left(\left\{x\in B_{\delta}\left(v\right):f(x)=f(v)\right\}\right)>\tau$]
-7:     $\Delta\leftarrow\Delta+[\delta_{\text{opt}}]$
-8:  end for
-9:  return  $\Delta$
-Using a sparse representation of the adjacency graph enables DCoM to handle large datasets efficiently without exhaustive space requirements.
-The algorithm’s complexity, including adjacency graph construction and sample selection, is discussed in App.
-C.
+Algorithm 1 DCoM, Active sampling
+Input: unlabeled pool $\mathbb{U}$}, labeled pool $\mathbb{L}$}, query size $q$, list of ball sizes $\Delta$, trained model $\mathbb{M}$} if $\mathbb{L}$$\neq\emptyset$
+Output: a set of points to query, and the coverage of $\mathbb{L}$}
+[…]
+2: Compute $P[C({\mathbb{L}},\Delta)]$
+3: Compute ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$
+[…]
+[…]
+6:for $(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta$ do
+7:  $\forall x\in B_{\delta_{i}}(x_{i})$: $\forall e=(v,x)\in E$, remove $e$
+8:  $\forall e=(x_{i},v)\in E$, remove $e$
+9:end for
+[…]
+11:for i=1,,qi=1,...,q do
+12:  $\forall x\in{\mathbb{U}}$, $\text{ODR}(x)\leftarrowOut$-Degree Rank
+13:  $\forall x\in{\mathbb{U}}$, $R(x)\leftarrow{S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))$ + (1 - ${S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x)$
+14:  $x_{max}\leftarrow\{\text{argmax}_{x\in{\mathbb{U}}}R(x)\}$
+15:  ${\textit{Q}}\leftarrow{{\textit{Q}}}\cup\{x_{max}\}$
+16:  $\forall x\in B_{\delta_{avg}}(x_{max})$: $\forall e=(v,x)\in E$, remove $e$
+17:  $M(x_{max})\leftarrow\text{1}$
+18:end for
+19:return Q, $P[C({\mathbb{L}},\Delta)]$
+Step 3: Model training: Obtain labels for the query set Q derived in Step 2.
+Remove Q from $\mathbb{U}$} and add it to $\mathbb{L}$}.
+Finally, train model $\mathbb{M}$} using the updated $\mathbb{L}$} (and also $\mathbb{U}$} if the learner employs semi-supervised learning).
+Step 4: $\boldsymbol{\Delta}$-Adjustment.
+Determine a suitable $\delta_{i}$ for each new labeled example $x_{i}\in{\textit{Q}}$ and add it to $\Delta$, see Alg 2 for pseudo-code and illustration in Fig.
+2.
+Begin by using the updated model $\mathbb{M}$} to predict labels for all points in the unlabeled set $\mathbb{U}$}.
+Likewise, compute a purity threshold $\tau=m\cdot P[C({\mathbb{L}}\setminus{\textit{Q}},\Delta)]+d$, where [^2] The linear transformation defined by $m,d$ adjusts the range of coverage scores into the relevant range of purity scores $[b,b+m]$.
+$P[C({\mathbb{L}}\setminus{\textit{Q}},\Delta)]$, the probability of coverage using the old labeled set, is derived in Step 2.
+Finally, for each new labeled example $v\in{\textit{Q}}$, search for the largest radius $\delta_{\text{opt}}$ whose purity $P\left(\left\{x\in B_{\delta_{\text{opt}}}\left(v\right):f(x)=f(v)\right\}\right)$ remains above $\tau$.
+Assuming that $\pi(B_{\delta}(x))$ is monotonic with $\delta$, use binary search.
+Algorithm 2 DCoM, $\boldsymbol{\Delta}$-Adjustment
+Input: unlabeled pool $\mathbb{U}$}, labeled pool $\mathbb{L}$}, query pool Q, maximal ball size $\delta_{\text{max}}$, list of ball sizes $\Delta$ for ${\mathbb{L}}\setminus{{\textit{Q}}}$, trained model $\mathbb{M}$} from the current iteration and $P[C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta)]$ from Step 2
+Output: updated list of ball sizes $\Delta$
+[…]
+[…]
+3:for $v\in{{\textit{Q}}}$ do
+4:  $\delta_{\text{opt}}\leftarrow\text{argmax}_{\delta}[P\left(\left\{x\in B_{\delta}\left(v\right):f(x)=f(v)\right\}\right)>\tau$]
+5:  $\Delta\leftarrow\Delta+[\delta_{\text{opt}}]$
+6:end for
+7:return $\Delta$
 4 Empirical evaluation
@@
 This section investigates two deep AL frameworks: a fully supervised approach, wherein a deep model is trained exclusively on the annotated dataset $\mathbb{L}$} as a standard supervised task, and a semi-supervised framework trained on both the annotated and unlabeled datasets, $\mathbb{L}$} and $\mathbb{U}$}.
-Our experimental setup is based on the codebase of [24], ensuring a fair comparison among the various AL strategies by using the same network architectures and sharing all relevant experimental conditions.
-More specifically, in the following experiments, we trained ResNet-18 [16] on STL-10 [6], SVHN [25], CIFAR-100 [22] and subsets of ImageNet [9] including ImageNet-50, ImageNet-100, and ImageNet-200 as they appear in [33].
-The hyper-parameters we used are detailed in App.
-A.
-In the semi-supervised framework, we used FlexMatch [42] and the code provided by [35].
-For the experiments, we followed the parameters specified by [42].
-The exact parameters are detailed in App.
-A.
+Our experimental setup is based on the codebase of (Munjal et al.
+2022), ensuring a fair comparison among the various AL strategies by using the same network architectures while sharing all the relevant experimental conditions.
+More specifically, we trained ResNet-18 (He et al.
+2016) on the following benchmark datasets: STL-10 (Coates, Ng, and Lee 2011), SVHN (Netzer et al.
+2011), CIFAR-100 (Krizhevsky, Hinton et al.
+2009) and subsets of ImageNet (Deng et al.
+2009) including ImageNet-50, ImageNet-100, and ImageNet-200 as defined in (Van Gansbeke et al.
+2020).
+The hyper-parameters are detailed in Appendix A.
+In the semi-supervised framework, we used FlexMatch (Zhang et al.
+2021) and the code provided by (Wang et al.
+2022), adopting the parameters specified by (Zhang et al.
+2021) (see details in Appendix A).
 While the ResNet-18 architecture may no longer achieve state-of-the-art results on the datasets examined, it serves as a suitable platform to evaluate the efficacy of AL strategies in a competitive and fair environment, where they have demonstrated benefits.
-In our comparisons, we used several active learning strategies (or optimization criteria) for selecting a query set: (1) Randomsampling, (2) Min margin— lowest margin between the two highest softmax outputs, (3) Max entropy— highest entropy of softmax outputs [30], (4) Uncertainty— 1 minus maximum softmax output, (5) DBAL[11], (6) CoreSet[28], (7) BALD[21], (8) BADGE[1], (9) ProbCover[39], and (10) LDM-s[4].
-When available, we used the code provided in [24] for each strategy.
+In our comparisons, we used several active learning strategies for selecting a query set: (i) Randomsampling, (ii) Min margin— lowest margin between the two highest softmax outputs; (iii) Max entropy— highest entropy of softmax outputs; (iv) Uncertainty— 1 minus the highest softmax output; (v) DBAL(Gal, Islam, and Ghahramani 2017); (vi) CoreSet(Sener and Savarese 2018); (vii) BALD(Kirsch, Van Amersfoort, and Gal 2019); (viii) BADGE(Ash et al.
+2019); (ix) ProbCover(Yehuda et al.
+2022); and (x) LDM-s(Cho et al.
+2023).
+When available, we used the code provided in (Munjal et al.
+2022) for each strategy.
 When it was unavailable, we used the code from the repositories of the corresponding papers.
-Since [4] doesn’t provide code, we compared our results with the corresponding available results reported in their paper, despite differences in the running setup (see App.
-D.6).
+Since (Cho et al.
+2023) doesn’t provide code, we compared our results with the corresponding available results reported in their paper, despite differences in the running setup (see Appendix E.5).
 DCoM and ProbCover require a suitable data embedding.
-For STL-10 , SVHN, CIFAR-10 and CIFAR-100 we employed the SimCLR [3] embedding.
-For the ImageNet subsets we used the DINOv2 [26] embedding.
+For STL-10 , SVHN, CIFAR-10 and CIFAR-100 we employed the SimCLR (Chen et al.
+2020) embedding.
+For the ImageNet subsets we used the DINOv2 (Oquab et al.
+2023) embedding.
 Section 4.3 provides an extensive ablation study, which includes alternative embedding spaces.
-In all experiments, identical settings and hyper-parameters are used for our method, as detailed in App.
-A.
-The robustness of our method to these choices is discussed in App.
-B.
+In all experiments, identical settings and hyper-parameters are used as detailed in Appendix A.
+The robustness of our method to these choices is discussed in Appendix C.
+Using a sparse representation of the adjacency graph enables DCoM to handle large datasets efficiently without exhaustive space requirements.
+The algorithm’s complexity, including adjacency graph construction and sample selection, is discussed in Appendix D.
 4.2 Main results
@@
 (f) ImageNet-200
-Figure 1: Mean accuracy difference between several AL algorithms and random queries.
+Figure 3: Mean accuracy difference between AL algorithms and random queries.
 Positive mean difference indicates that the corresponding AL method is beneficial.
-The errors bars correspond to standard error, based on 5-10 repetitions (3 for the ImageNet subsets).
-While some methods perform well only in specific budget regimes, DCoM consistently achieves the best results across all budgets.
+The errors bars correspond to standard error, using 5-10 repetitions (3 for the ImageNet subsets).
+While some methods perform well only in specific budget regime, DCoM consistently achieves the best results across all budgets.
+Cumulative budget is evenly spaced for easy range comparison.
 Fully supervised framework.
@@
 3 shows the difference between the accuracy obtained by each AL method, and the accuracy obtained when training a similar network with a random query set of the same size (the baseline of no active learning), for each AL iteration.
-The final accuracy of each method is shown in Tables 1-2 for the ImageNet-100 and CIFAR-100 datasets, as mean $\pm$ Standard Error (STE).
-Additional datasets are shown in App.
-D.
-Table 1: Model final accuracy when varying the size of the labeled set ${\mathbb{L}}$ (row) and the respective AL strategy (column), using the ImageNet-100 dataset
-$|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin DCoM
-200 4.58±0.13 6.72±0.07 4.67±0.16 1.99±0.22 4.58±0.13 2.01±0.06 1.99±0.10 1.83±0.06 1.84±0.12 6.87±0.18
-400 6.61±0.17 9.30±0.23 6.42±0.23 4.39±0.39 5.53±0.16 3.50±0.32 4.17±0.33 3.65±0.25 4.71±0.10 10.59±0.26
-600 8.11±0.42 12.71±0.42 8.01±0.20 6.42±0.24 6.45±0.37 4.93±0.29 5.51±0.33 4.96±0.22 6.46±0.12 13.61±0.15
-1,000 11.43±0.17 17.71±0.27 11.56±0.38 9.42±0.19 9.04±0.66 7.97±0.82 8.39±0.18 7.45±0.18 10.43±0.11 17.53±0.23
-3,000 27.59±0.76 34.01±0.15 27.58±0.72 24.96±0.23 19.41±0.32 23.10±0.60 20.97±0.33 20.88±0.52 26.51±0.79 34.85±0.23
-5,000 39.25±0.26 41.91±0.20 39.13±0.06 36.50±0.25 29.35±0.39 33.78±0.46 31.89±0.24 31.25±0.21 38.37±0.10 44.09±0.83
-7,000 46.88±0.25 48.41±0.42 46.42±0.41 43.49±0.42 38.45±0.12 42.09±0.54 41.19±0.17 40.47±0.83 45.69±0.26 49.22±0.29
-11,000 55.31±0.26 54.01±0.15 56.61±0.11 55.29±0.32 50.63±0.33 53.25±0.31 52.52±0.26 53.09±0.29 56.86±0.10 58.45±0.15
-15,000 61.69±0.09 57.04±0.43 63.05±0.42 62.19±0.21 58.71±0.05 60.07±0.07 59.76±0.22 59.64±0.30 62.71±0.27 64.57±0.22
-19,000 66.39±0.26 59.69±0.16 66.82±0.33 66.19±0.19 64.77±0.32 65.59±0.27 64.65±0.23 64.93±0.23 66.89±0.28 67.60±0.08
-23,000 68.92±0.15 61.99±0.41 70.54±0.21 69.62±0.32 68.49±0.06 69.03±0.20 68.47±0.50 67.65±0.14 70.81±0.35 70.21±0.32
-Table 2: Same as Table 1, using the CIFAR-100 dataset.
+The final accuracy of each method is shown in Table 1 for the CIFAR-100 dataset, as mean $\pm$ Standard Error (STE).
+Similar tables for the remaining datasets are shown in Appendix E.
 $|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin LDM-s DCoM
@@
 23,000 57.74±0.14 56.29±0.21 59.60±0.16 59.96±0.35 59.59±0.22 59.79±0.23 59.91±0.08 59.72±0.21 59.88±0.21 52.48±0.00 59.60±0.33
+Table 1: Model accuracy with varying labeled set sizes $\mathbb{L}$} (row) and AL strategies (column) while training directly on CIFAR-100 images.
 Semi-supervised framework.
@@
 Figure 4: Comparison of AL strategies in a semi-supervised learning framework.
-Each bar represents the mean test accuracy over 3 repetitions of FlexMatch, trained using a total of 400 (left) and 1,000 (right) labeled examples on the CIFAR-100 dataset (an average of 4 and 10 labeled examples per class respectively).
-Three rounds of AL are considered, selecting 100 examples in the first round, 300 in the second, and 600 in the third.
+Each bar shows the mean test accuracy over 3 repetitions of FlexMatch, trained with 400 and 1,000 labeled examples on CIFAR-100.
+Three AL rounds are conducted, selecting 100, 300, and 600 examples in each round.
 4.3 Ablation study
@@
 The results demonstrate the robustness of our method to these choices.
-Additional choices are investigate in App.
-D, including the contribution of a dynamic $\delta$ and initialization $\delta_{0}$ (App.
-D.1).
-In App.
-D.3, we evaluate the effectiveness of ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ from (6) in capturing competence.
-Different feature spaces.
-As discussed in Section 3.1, our approach relies on the existence of a suitable embedding space, where distance is indicative of semantic dissimilarity.
-We now repeat the basic fully-supervised experiments while varying the embedding employed by DCoM, considering various popular self-supervised representations.
-Results are reported in Fig.
-4, showing that DCoM consistently delivers the best performance regardless of the specific embedding used.
-Breaking down the algorithm into its components.
-We evaluate the contribution of each component in objective function (1) separately.
-Results are shown in Fig.
-4, wherein DPC denotes the Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM, and Margin denotes DCoM’s choice for ${{\mathcal{O}_{high}}}(x)$.
-Clearly, the weighted objective function of DCoM yields superior results.
-Similar behavior is observed for several other datasets and other selections of ${{\mathcal{O}_{high}}}(x)$, see App.
-D.4.
+Additional choices are investigate in Appendix E, including the contribution of a dynamic $\delta$ and initialization $\delta_{0}$ (Appendix E.1).
+In Appendix E.3, we evaluate the effectiveness of ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ from (5) in capturing competence.
 (a) MOCOv2+
@@
 Figure 5: Performance over CIFAR-100, using two different embedding spaces for DCoM and ProbCover, see details in the caption of Fig.
-1.
+3.
 Clearly, DCoM consistently achieves the best results.
-Results using another embedding space are provided in App.
-D.5.
-(c) CIFAR-10
-Figure 4: Performance evaluation (mean accuracy and STE) when optimizing each component of objective function (1) separately and together.
+Results using an additional embedding space are shown in Appendix E.4.
+Different feature spaces.
+As discussed in Section 3, our approach relies on the existence of a suitable embedding space, where distance is indicative of semantic dissimilarity.
+We now repeat the basic fully-supervised experiments while varying the embedding space, comparing various popular self-supervised representations.
+Results are reported in Fig.
+5, showing that DCoM consistently delivers the best performance regardless of the specific embedding used.
+Breaking down the algorithm into its components.
+The objective function in (1) includes the linear combination of two objective functions, weighed by competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ from (5).
+We evaluate the contribution of each component in this objective function separately.
+Results are shown in Fig.
+6, wherein DPC denotes the Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM, and Margin denotes DCoM’s choice for ${{\mathcal{O}_{high}}}(x)$.
+Clearly, the weighted objective function of DCoM yields superior results.
+(a) CIFAR-10
+(b) IMAGENET-50
+Figure 6: Performance (mean accuracy and STE) when optimizing each component of objective function (1) separately and together.
 Different uncertainty-based methods.
-We now evaluate our choice for ${{{\mathcal{O}_{high}}}(x)}$ in objective function (1), investigating the alternative methods of min margin, entropy of the min max activation, and uncertainty (see details in Section 4.1).
-Additionally, we use a score reminiscent of the AL method BADGE [11] and termed Gradient norm, which computes the norm of the gradient between the learner’s last two layers.
+We evaluate our choice for ${{\mathcal{O}_{high}}}(x)$} in objective function (1), investigating the alternative methods of min margin, entropy of the min max activation, and uncertainty (see details in Section 4.1).
+Additionally, we use a score reminiscent of the AL method BADGE (Gal, Islam, and Ghahramani 2017) and termed Gradient norm, which computes the norm of the gradient between the learner’s last two layers.
 Once again, the results of DCoM show robustness to this choice, with insignificant differences between the different uncertainty scores, see Fig.
-5.
+7.
 Consequently, we prioritize computational efficiency and opt for the fastest-to-compute score, which is min-margin.
-Table 3: Model accuracy for different sizes of labeled set ${\mathbb{L}}$ and several uncertainty-based methods for ${\mathbb{F}}(x)$
-$|{\mathbb{L}}|$ Random Entropy Margin Uncertainty Gradient norm
-200 6.39±0.19 10.71±0.30 10.59±0.19 10.44±0.26 10.33±0.21
-400 8.74±0.07 14.35±0.24 14.11±0.11 14.72±0.20 14.59±0.19
-600 10.73±0.13 17.07±0.19 16.66±0.11 16.75±0.34 16.83±0.17
-1,000 13.49±0.18 19.98±0.22 20.21±0.07 20.40±0.21 20.03±0.17
-3,000 24.02±0.27 29.89±0.11 30.36±0.08 30.30±0.11 30.62±0.52
-5,000 31.52±0.39 36.08±0.13 36.95±0.10 36.06±0.18 36.12±0.46
-7,000 37.69±0.16 41.18±0.25 41.52±0.25 41.80±0.08 41.23±0.65
-11,000 45.66±0.22 47.23±0.09 46.86±0.28 46.99±0.27 46.84±0.28
-15,000 50.57±0.09 52.67±0.15 52.33±0.22 52.43±0.14 52.21±0.22
-19,000 54.95±0.09 56.64±0.27 56.53±0.07 56.47±0.06 56.35±0.28
-23,000 57.74±0.14 59.50±0.18 59.60±0.33 59.68±0.24 59.45±0.06
 Figure 7: Similar to Fig.
-1 and Tables 1-2, using only DCoM while leveraging 4 different variants for its underlying uncertainty score (see text).
-The performance differences between the variants are small.
+3, using only DCoM while leveraging 4 different variants for its underlying uncertainty score (see text).
+The performance differences between the variants are hardly significant.
 5 Summary
@@
 The results show that DCoM significantly outperforms alternative methods across all budgetary constraints.
+Acknowledgments
+This work was supported by grants from the Israeli Council of Higher Education, AFOSR award FA8655-24-1-7006, and the Gatsby Charitable Foundation.
 Appendix
 Appendix A Implementation details
-The source code used in this study is available at https://github.com/avihu111/TypiClust.
-The code is based on the GitHub repositories of Yehuda et al. [39] and Munjal et al. [24].
+The source code for this study can be found in the zip file included in the supplementary material.
+This zip file contains a USAGE.md file that provides instructions on how to run the algorithm.
+The code in the zip file is based on the GitHub repositories of Yehuda et al.
+(2022) and Munjal et al.
+(2022).
+Our GitHub repository will be made public upon acceptance.
 DCoM necessitates the computation of the initial delta for each dataset representation.
 The values utilized in our experiments can be found in Table 2.
-The $\delta$ values of CIFAR-10 and CIFAR-100 are sourced from [39].
-Information on the computation of $\delta$ values for other datasets can be located in App.
-D.1.
-Table 4: Initial delta values used in experiments
-Dataset SSL method Initial $\delta$
+The $\delta$ values of CIFAR-10 and CIFAR-100 are sourced from (Yehuda et al.
+2022).
+Information on the computation of $\delta$ values for other datasets can be located in Appendix E.1.
+Dataset SSL method $\delta_{0}$
 STL-10 SimCLR 0.55
@@
 CIFAR-100 Barlow Twins 0.55
+Table 2: Initial delta values used in experiments
 Additionally, throughout our experiments we used identical algorithm parameters, distinguishing between those for datasets with smaller class amounts (10) and larger ones (50+).
@@
 The $\delta$ resolution for the binary search in DCoM was set to $0.05$.
-Our ablation study, detailed in App.
-B, demonstrates that these selections have negligible impact on performance.
+Our ablation study, detailed in Appendix C, demonstrates that these selections have negligible impact on performance.
 A.1 Supervised training
@@
 Training was performed with a batch size of 100 examples, and we applied random cropping and horizontal flips for data augmentation.
-For an illustration of these parameters in use, refer to [24].
+For an illustration of these parameters in use, refer to (Munjal et al.
+2022).
 When training ImageNet-50, we used the same hyper-parameters, only changing the base learning rate to $0.01$, the batch size to $50$ and the epoch amount to $50$.
@@
 A.2 Semi-supervised training
-When training FlexMatch [42], we used the semi supervised framework by [35], and the AL framework by [24]. All experiments involved 3 repetitions.
-We relied on the standard hyper-parameters used by FlexMatch [42].
+When training FlexMatch (Zhang et al. 2021), we used the semi supervised framework by (Wang et al. 2022), and the AL framework by (Munjal et al. 2022). All experiments involved 3 repetitions.
+We relied on the standard hyper-parameters used by FlexMatch (Zhang et al.
+2021).
 Specifically, we trained WideResNet-28 for 512 epochs using the SGD optimizer, with $0.03$ learning rate, $64$ batch size, $0.9$ SGD momentum, $0.999$ EMA momentum, $0.001$ weight decay and $2$ widen factor.
@@
 STL-10, SVHN, CIFAR-10, CIFAR-100.
-To extract semantically meaningful features, we trained SimCLR using the code provided by [33] for STL-10, SVHN, CIFAR-10 and CIFAR-100.
-Specifically, we used ResNet-18 [16] with an MLP projection layer to a $128$-dim vector, trained for 512 epochs.
-All the training hyper-parameters were identical to those used by SCAN (all details can be found in [33]).
+To extract semantically meaningful features, we trained SimCLR using the code provided by (Van Gansbeke et al.
+2020) for STL-10, SVHN, CIFAR-10 and CIFAR-100.
+Specifically, we used ResNet-18 (He et al.
+2016) with an MLP projection layer to a $128$-dim vector, trained for 512 epochs.
+All the training hyper-parameters were identical to those used by SCAN (all details can be found in (Van Gansbeke et al.
+2020)).
 After training, we used the $512$ dimensional penultimate layer as the representation space.
 Representation learning: ImageNet-50/100/200.
-We extracted features from the official (ViT-S/14) DINOv2 weights pre-trained on ImageNet [26].
+We extracted features from the official (ViT-S/14) DINOv2 weights pre-trained on ImageNet (Oquab et al.
+2023).
 We employed the L2-normalized penultimate layer for the embedding, which has a dimensionality of $384$.
-Ablation embeddings – CIFAR-100. We extracted more features as BYOL [13], MOCOv2+ [17], and Barlow Twins [41] using pre-trained weights from [8].
-Appendix B Hyper-parameters exploring
-As described before, DCoM has 3 hyper-parameters: $a,k\text{and }\tau$.
+Ablation embeddings – CIFAR-100.
+We extracted more features as BYOL (Grill et al.
+2020), MOCOv2+ (He et al.
+2020), and Barlow Twins (Zbontar et al.
+2021) using pre-trained weights from (da Costa et al.
+2022).
+Appendix B Purity Calculation in nNN
+When transitioning from 1NN to nNN, the structure of the problem remains unchanged, with the key difference being that the radius $\delta$ is adjusted for each labeled example.
+Next, we demonstrate that the purity of each ball in the nNN problem can be precisely computed using these $\delta$-values.
+This is because when a labeled example $c\in{\mathbb{L}}$ assigns a label to a point $x$, it must also cover $x$.
+Therefore, any mislabeling by $c$ would reduce the purity, as illustrated by the following statement:
+Proposition.
+If $c\in{\mathbb{L}}$ is labeling $x$ and $x\in C({\mathbb{L}},\Delta)$, then $x\in B_{\delta_{c}}(c)$.
+Proof.
+Let $c\in{\mathbb{L}}$ be the nearest neighbor (nNN) of $x$, and let $\delta_{c}$ be the corresponding radius for $c$.
+By definition, $f^{nNN}(x)=f(x)$, and since $c$ is labeled, we have $f(c)=f^{nNN}(c)\equiv f^{N}(c)$.
+Given that $x\in C({\mathbb{L}},\Delta)$, there exists a labeled example $c^{\prime}\in{\mathbb{L}}$ that covers $x$, meaning $d(x,c^{\prime})<\delta_{c^{\prime}}$.
+Suppose that $c\neq c^{\prime}$.
+Since $c$ is the nNN of $x$, it follows that:
+$\frac{d(x,c)}{\delta_{c}}<\frac{d(x,c^{\prime})}{\delta_{c^{\prime}}}$
+Given $d(x,c^{\prime})<\delta_{c^{\prime}}$, we get:
+$\frac{d(x,c)}{\delta_{c}}<1\Rightarrow d(x,c)<\delta_{c}$
+Thus, $x\in B_{\delta_{c}}(c)$, meaning $c$ covers $x$ within the radius $\delta_{c}$.
+This implies that if $x$ were wrongly labeled, $c$ would still cover $x$, indicating lower purity in $B_{\delta_{c}}(c)$. ∎
+Appendix C Hyper-parameters exploring
+Theoretical basis for the Choice of $\boldsymbol{a}$ and $\boldsymbol{k}$.
+The parameter $a$ represents the center of the sigmoid function.
+Our objective is for the algorithm to assign increased weight to $\mathcal{O}_{high}$} only after it has covered a substantial portion of the dataset.
+Consequently, we choose $a\in(0.5,1)$.
+The parameter $k$ controls the steepness of the sigmoid curve.
+The function must yield 0 in the absence of coverage and approach 1 as coverage becomes complete.
+To achieve this, we configure the numerator of the fraction to be equal to the denominator when the cover’s probability is 1.
+Although the function cannot pass precisely through $(0,0)$, it can asymptotically approach $(0,\varepsilon)$ for an infinitesimally small $\varepsilon\in\Re$.
+We now investigate the relationship between this approximation error and the parameter $k$. Define the function $f(x)$ as $S_{\mathcal{L}}$} for simplicity:
+$f(x)=\frac{1+e^{-k(1-a)}}{1+e^{-k(x-a)}}$
+The limit of $f(x)$ as $x$ approaches 0 is
+$\lim_{x\to 0}f(x)=\frac{1+e^{-k(1-a)}}{1+e^{ka}}$
+For the function to approach 0, we require:
+$\frac{1+e^{-k(1-a)}}{1+e^{ka}}\ll 1$
+This implies:
+$1+e^{-k(1-a)}\ll 1+e^{ka}$
+$e^{-k(1-a)}\ll e^{ka}$
+$e^{-k}\ll e^{ka}$
+$e^{-k}\ll 1$
+Next, we calculate $f(0)$:
+$f(0)=\frac{1+e^{-k(1-a)}}{1+e^{ka}}=\varepsilon$
+Solving for $k$:
+$1+e^{-k(1-a)}=\varepsilon(1+e^{ka})$
+$1-\varepsilon=\varepsilon e^{ka}-e^{-k(1-a)}$
+$1-\varepsilon=e^{ka}(\varepsilon-e^{-k})$
+$e^{ka}=\frac{1-\varepsilon}{\varepsilon-e^{-k}}$
+Approximating $e^{ka}$ using (*):
+$e^{ka}\approx\frac{1-\varepsilon}{\varepsilon}$
+So:
+$ka\approx\ln\left(\frac{1-\varepsilon}{\varepsilon}\right)$
+Finally:
+$k\approx\frac{1}{a}\ln\left(\frac{1}{\varepsilon}-1\right)\approx\frac{\ln\left(\frac{1}{\varepsilon}\right)}{a}$
+To minimize the approximation error, we should choose a larger $k$. In our experiments with $a=0.8\text{ or }0.9$ and $\varepsilon=10^{-10}$, we determined:
+$k(a=0.9,\varepsilon=10^{-10})\approx 25.58$
+$k(a=0.8,\varepsilon=10^{-10})\approx 28.78$
+The value of $\varepsilon$ was chosen to be asymptotically smaller than the dimension of the initial graphs constructed for all the datasets. In accordance, we chose $k=30$.
+Experimental assessment of DCoM hyper-parameter consistency As described before, DCoM has 3 hyper-parameters: $a,k\text{ and }\tau$.
 When running the main experiments described in Section 4.2, we maintained consistency by employing identical parameters for STL-10, SVHN, and CIFAR-10, as well as for CIFAR-100 and the subsets of ImageNet.
@@
 Here, we conduct experiments with DCoM using STL-10 and CIFAR-100 datasets and several values for $a$ and $k$ from the definition of the competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ in (5).
-We observe that while certain choices may be slightly better than others, all choices yield results that are very close to each other, and DCoM consistently outperforms or matches the performance of previous methods.
+We observe that while certain choices may be slightly better than others, all choices yield similar results, and DCoM consistently outperforms or matches the performance of previous methods.
 (a) STL-10
@@
 Figure 8: Evaluation of DCoM across different settings of the logistic function hyper-parameters a and k.
-Each plot displays the mean and standard error over $3-5$ repetitions.
+Each plot displays the mean and standard error over 353-5 repetitions.
 The findings indicate that fine-tuning these parameters has negligible effects on performance.
-Appendix C Time and space complexity of DCoM
-For the complexity analysis below, let $n$ represent the number of examples in the combined unlabeled and labeled pool $|{\mathbb{L}}\cup{\mathbb{U}}|$, $d$ the dimension of the data embedding space, $b$ the given budget ($|{\mathbb{L}}|$) and $q$ the size of the query set.
+Appendix D Time and space complexity of DCoM
+For the complexity analysis below, let $n$ represent the number of examples in the combined unlabeled and labeled pool $|{\mathbb{U}}\cup{\mathbb{L}}|$, $d$ the dimension of the data embedding space, $b$ - the given budget ($|{\mathbb{L}}|$) and $q$ - the size of the query set.
 As previously stated, DCoM can be divided into three distinct steps:
-C.1 Selection of initial $\boldsymbol{\delta}$
+D.1 Selection of initial $\boldsymbol{\delta}$ delta
 Time Complexity: The process begins by creating $t$ adjacency graphs, each corresponding to a different $\delta$ value being evaluated.
@@
 The computation of purity requires the prediction of pseudo-labels by k-means at a cost of $O(n^{2})$.
-The complexity of computing ball purity for each example is $O(n^{2})$, so totally each the ball purity computing takes $O(n^{3})$.
-This adds up to overall time complexity of $O(tdn^{3})$.
+The complexity of computing ball purity for one example is $O(n^{2})$, so totally the ball purity computing for one graph takes $O(n^{3})$.
+This adds up to overall time complexity of $O(tdn^{2}+n^{2}+tn^{3})=O(tn^{3})$.
+We can construct the initial graph using the largest $\delta$-value and then, instead of building a new graph for each $\delta$-value in $O(tdn^{2})$, we can filter distances above the current $\delta$ in $O(dn^{2}+n^{2})$.
+However, the total complexity remains $O(tn^{3})$.
 Although it may seem significant, this step doesn’t occur during the active learning algorithm process.
-Additionally, it’s possible to confine the data employed for purity computation to the densest points, effectively lowering the complexity to $O(tdn^{2})$.
+Additionally, it’s possible to confine the data employed for purity computation to the densest points, effectively lowering the complexity to $O(dn^{2})$.
 This approach mandates the use of at least the number of points equivalent to the number of classes for purity computation.
@@
 Space complexity: Naively, the space complexity is $O(n^{2})$, which might be impractical for large datasets like ImageNet.
-However, [39] demonstrates that utilizing a sparse matrix in coordinate list (COO) format and using limited $\delta$ values, the space complexity becomes $O(|E|)$, where $E$ represents the set of edges in the graph.
+However, (Yehuda et al.
+2022) demonstrates that utilizing a sparse matrix in coordinate list (COO) format and using limited $\delta$ values, the space complexity becomes $O(|E|)$, where $E$ represents the set of edges in the graph.
 Although $O(|E|)$ remains $O(n^{2})$ in the worst-case scenario, in practice, the average vertex degree with a limited radius is less than $n$.
-C.2 Customizing $\boldsymbol{\delta}$ values
+D.2 $\boldsymbol{\Delta}$-Adjustment
 Time complexity: Each iteration of DCoM starts with customizing $\delta$ values for each labeled example from the last active step.
@@
 This implies that there are $\frac{\delta_{\text{max}}}{r}$ available values for $\delta$.
-Overall, this process takes $O(dn^{2}+n+|{\mathbb{L}}|\cdot\log(\frac{\delta_{\text{max}}}{r}))=O(dn^{2}+b\cdot\log(\frac{\delta_{\text{max}}}{r}))$.
+For each $\delta$ value, the algorithm computes its purity value.
+Overall, this process takes $O(dn^{2}+n+|{\mathbb{L}}|\cdot\log(\frac{\delta_{\text{max}}}{r})\cdot n^{2})=O(dn^{2}+b\cdot\log(\frac{\delta_{\text{max}}}{r})\cdot n^{2})$.
 In practical terms, due to the vectorization of these processes, it requires approximately 27 minutes to construct an adjacency graph for ImageNet-200 and update deltas for 1000 examples on a single CPU.
-Space complexity: Similar to the space complexity analysis in App. C.1, which is $O(n^{2})$.
-C.3 Active sampling
+Space complexity: Similar to the space complexity analysis in Appendix D.1, which is $O(n^{2})$.
+D.3 Active sampling
 Time complexity: Active sampling involves balancing model min-margin and example density based on the current adjacency graph.
@@
 Subsequently, the current adjacency is created using $\delta_{\text{avg}}$, which is the average over the $\mathbb{L}$} $\delta$ values list - $\Delta$.
-As discussed in App.
-C.1, this step has a time complexity of $O(dn^{2})$.
+As discussed in Appendix D.1, this step has a time complexity of $O(dn^{2})$.
 Following this, the iterative process for selecting a single sample includes the following steps:
@@
 Space complexity: Similar to the previous space complexity analysis – $O(n^{2})$.
-Appendix D Additional empirical results
-D.1 $\delta_{0}\,\,\,$ Initialization
-To set the initial value $\delta_{0}$, we adopted the method outlined in [39], as detailed in Section 3.
+Appendix E Additional empirical results
+E.1 $\boldsymbol{\delta_{0}}$ Initialization
+To set the initial value $\delta_{0}$, we adopted the method outlined in (Yehuda et al.
+2022), as detailed in Section 3.
 Fig.
 9 displays the purity function across various $\delta$ values, along with the selected $\delta$ values for each dataset.
-As mentioned previously, the $\delta$ values for CIFAR-10 and CIFAR-100 are derived from [39].
+As mentioned previously, the $\delta$ values for CIFAR-10 and CIFAR-100 are derived from (Yehuda et al.
+2022).
 (a) STL-10
@@
 (d) ImageNet-100
-(e) ImageNet-200
 Figure 9: $\pi(\delta)$, estimated from the unlabeled data and using k-means algorithm for labeling. The dashed line indicates the highest $\delta$, after which purity drops below $\alpha=0.95$.
-D.2 $\Delta\,\,\,\,$ distribution throughout the training process
+E.2 $\boldsymbol{\Delta}$ distribution evolution throughout training
 The experiment in Fig.
@@
 (b) CIFAR-100
-(c) ImageNet-50
 Figure 10: The $\Delta$ distribution through DCoM algorithm.
-Each plot displays the mean and standard error over 5-10 repetitions.
+Each plot displays the distribution over 5-10 repetitions.
 The results indicate that as the emphasis shifts towards maximizing the purity of sample balls, rather than solely focusing on maximum coverage, the radii adapt accordingly.
 This, combined with the improved performance observed with DCoM, highlights the necessity of the dynamic algorithm in enhancing the effectiveness of active sampling.
-D.3 Dynamic algorithm coverage
-The analysis presented in Fig.
-9 demonstrates a consistent behavior of probability coverage across different datasets during the DCoM process.
-This indicates that using probability coverage as a measure of progress consistently proves to be a wise decision.
-(a) CIFAR-10 and STL-10
-(b) CIFAR-100 and ImageNet-50
-Figure 9: Comparison between $P(C({\mathbb{L}},\Delta))$ values during DCoM sampling process, across several datasets.
-Each plot displays the mean and standard error over $3-10$ repetitions.
-These figures demonstrates a consistent behavior of probability coverage.
-D.4 Decomposing the algorithm
-In Section 4.3, we present an ablation study where we evaluate the contribution of each component in objective function (1) separately on CIFAR-100 dataset.
-In Fig.
-10 you can see the results over CIFAR-100 and ImageNet-50 datasets.
-(a) CIFAR-100
-(b) ImageNet-50
-Figure 10: Performance evaluation (mean accuracy and STE) when optimizing each component of objective function (1) separately and together.
-DPC denotes the Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM.
-Margin is DCoM’s choice for ${{\mathcal{O}_{high}}}(x)$.
-In Fig. 11 you can see the same ablation but using entropy as ${{\mathcal{O}_{high}}}(x)$ instead of margin.
-(a) CIFAR-10
-(b) CIFAR-100
-(c) ImageNet-50
-Figure 11: Assessment of each algorithm’s objective function individually across multiple datasets.
-DPC stands for Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM.
-DCEM represents DCoM with entropy as ${{\mathcal{O}_{high}}}(x)$ instead of margin.
-Each plot displays the mean and standard error over $5-10$ repetitions.
-These graphs demonstrate how the weighted objective function consistently produces superior results across all labeled set sizes $|{\mathbb{L}}|$.
-D.5 Different embedding space
-In Section 4.3, we present an ablation study where we repeat the basic fully-supervised experiments while varying the embedding employed by DCoM.
-In Fig.
-12 you can see the same ablation using BYOL [13] representation.
-Figure 12: Performance over CIFAR-100, using BYOL feature space over DCoM and ProbCover, see details in the caption of Fig. 1. Clearly, DCoM consistently achieves the best results.
-D.6 Comparison per dataset
-Tables 5-6 present the empirical accuracy values of several datasets using different active learning algorithms.
-In each active step, a new learner is trained from scratch over all available labeled data, and the results are presented as the [mean $\pm$ STE] over $5-10$ repetitions (3 for ImageNet subsets).
-As mentioned earlier, the results for LDM-s algorithm were acquired from the study by [4].
-Their lack of code provision suggests discrepancies in the running setup.
-Nonetheless, the disparity between their method and random selection in their setup is smaller than that observed in ours, implying the superiority of our method.
-The table includes the results for all datasets, with the name of each dataset below it.
-Table 5: Model accuracy for different sizes of labeled set ${\mathbb{L}}$ and AL strategies
 $|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin LDM-s DCoM
@@
 ImageNet-100 dataset
+Table 3: Model accuracy for different sizes of labeled set $\mathbb{L}$} and AL strategies.
 $|{\mathbb{L}}|$ Random Prob Cover BALD Coreset Uncertainty Entropy DBAL Margin DCoM
@@
 ImageNet-200 dataset
-Table 6: Model accuracy for different sizes of labeled set ${\mathbb{L}}$ and AL strategies
 $|{\mathbb{L}}|$ Random Prob Cover BADGE BALD Coreset Uncertainty Entropy DBAL Margin DCoM
@@
 SVHN dataset
+Table 4: Model accuracy for different sizes of labeled set $\mathbb{L}$} and AL strategies.
+[…]
+Fig.
+11 tracks the behavior of the competence score $S_{\mathcal{L}}$} across 3 datasets.
+In all cases, the comparison of Figs.11(a),11(b) reveals that $S_{\mathcal{L}}$} increases monotonically with competence as measured by accuracy.
+Interestingly, while the critical point where $S_{\mathcal{L}}$} changes from favoring $\mathcal{O}_{low}$} (small values) to favoring $\mathcal{O}_{high}$} (large values) corresponds to an intermediate level of accuracy, this transition is delayed for the more difficult datasets.
+(a) Accuracy with random sampling.
+(b) $S_{\mathcal{L}}$} during DCoM training.
+Figure 11: (a) Accuracy (mean and standard error) as a function of budget $b$, learning to classify CIFAR-100, ImageNet-50, and ImageNet-100 using a random sample of $b$ points.
+(b) The competence score $S_{\mathcal{L}}$} as a function of budget.
+E.4 Different embedding space
+In Section 4.3, we present an ablation study where we repeat the basic fully-supervised experiments while varying the embedding employed by DCoM.
+In Fig.
+12 you can see the same ablation using BYOL (Grill et al.
+2020) representation.
+Figure 12: Performance over CIFAR-100, using BYOL feature space over DCoM and ProbCover, see details in the caption of Fig. 3. Clearly, DCoM consistently achieves the best results.
+E.5 Comparison per dataset
+Tables 3-4 present the empirical accuracy values of several datasets using different active learning algorithms.
+In each active step, a new learner is trained from scratch over all available labeled data, and the results are presented as the [mean $\pm$ STE] over 5105-10 repetitions (3 for ImageNet subsets).
+As mentioned earlier, the results for LDM-s algorithm were acquired from the study by (Cho et al.
+2023).
+Their lack of code provision suggests discrepancies in the running setup.
+Nonetheless, the disparity between their method and random selection in their setup is smaller than that observed in ours, implying the superiority of our method.
+The table includes the results for all datasets, with the name of each dataset below it.